### v1

# A Generalized Sylvester-Fermat-Torricelli problem with application in disaster relief operations by UAVs

Abstract
Natural and human-made disasters can cause severe devastation and claim thousands of lives worldwide. Therefore, developing efficient methods for disaster response and management is a critical task for relief teams. One of the most essential components of effective response is the rapid collection of information about affected areas, damages, and victims. More data translates into better coordination, faster rescue operations, and ultimately, more lives saved. However, in some disasters, such as earthquakes, the communication infrastructure is often partially or completely destroyed, making it extremely difficult for victims to send distress signals and for rescue teams to locate and assist them in time.
Unmanned Aerial Vehicles (UAVs) have emerged as valuable tools in such scenarios. In particular, a fleet of UAVs can be dispatched from a mobile station to the affected area to facilitate data collection and establish temporary communication networks. Nevertheless, real-world deployment of UAVs faces several challenges, with adverse weather conditions—especially wind—being among the most significant. To address this, we develop a novel mathematical framework to determine the optimal location of a mobile UAV station while explicitly accounting for the heterogeneity of the UAVs and the effect of wind. Our approach extends a classical single-facility location problem by incorporating heterogeneous dynamic sets that represent varying UAV speed. In particular, we generalize the Sylvester problem to introduce the Sylvester-Fermat-Torricelli (SFT) problem, which captures complex factors such as wind influence, UAV heterogeneity, and back-and-forth motion within a unified framework. The proposed framework enhances the practicality of UAV-based disaster response planning by accounting for real-world factors such as wind and UAV heterogeneity. Experimental results demonstrate that it can reduce wasted operational time by up to 84%, making post-disaster missions significantly more efficient and effective.
keywords:
Generalized Sylvester problem, Unmanned aerial vehicle (UAV), Generalized Sylvester-Fermat-Torricelli problem, Nonsmooth optimization, Convex analysis.
1 Introduction
1.1 Introduction and Motivation
Natural and human-made disasters worldwide can have devastating consequences, leading to significant loss of life and extensive destruction, often resulting in substantial economic and social costs. For instance, in 2023, the Emergency Events Database (EM-DAT) recorded 399 disasters worldwide, leading to 86,473 fatalities and affecting approximately 93.1 million people [10]. Given the significant costs associated with disasters, disaster management has become a critical field focused on minimizing these costs. Effective disaster management strategies can be categorized into pre-disaster activities (prevention and preparedness) and post-disaster (response and recovery) ones [8].
In post-disaster actions, one of the most critical tasks is data collection. This data includes the locations of victims trapped under debris, the extent of road destruction and accessibility, and the damage levels across different parts of the region of interest. In this context, UAVs offer significant advantages, including rapid area surveying, ease of deployment, low-cost maintenance, and the ability to reach inaccessible locations [18], making them highly effective for data gathering. Moreover, since natural disasters can damage or render network infrastructure inoperable [8], UAVs can also serve as a vital tool for establishing temporary communication networks. Therefore, several studies have explored the use of UAVs to provide wireless or Delay-Tolerant Networks (DTNs) in such scenarios [40, 20, 1, 8, 14, 9, 42, 43].
While UAVs can be valuable for providing network coverage, their main drawback is limited energy capacity, requiring them to return periodically for recharging [40]. Also, unlike other vehicles, UAVs are susceptible to wind due to their lighter weight, smaller size, lower flight altitude, and lower speed [11]. These limitations create significant challenges in deploying UAVs for post-disaster network provisioning. [8] proposed a mobile first-response UAV station with a fleet of heterogenous UAVs that has a long-distance communication antenna, an electricity generator and a system for automatic UAV battery recharging. Motivated by this idea, this paper aims to develop a reliable mathematical solution for determining the optimal location of a mobile UAV station, considering UAV heterogeneity and the impact of wind, which is common in real-world scenarios.
This problem can be regarded as a single-facility location problem, previously studied in the context of drone-assisted delivery, where drones launch from a truck, deliver parcels, and subsequently return to the truck. There are studies [26, 3, 38, 2] in the literature that examines the location of launch points (location of truck) for drones in a delivery network where a truck and multiple drones operate in tandem. [39] considered a tandem delivery strategy like [26] to use in the healthcare application. However, To the best of our knowledge, little attention has been given to addressing the impact of wind on drone performance. This factor cannot be ignored as it causes delays in delivery time and increases energy consumption. We also tried to consider the effect of wind in other applications that are search and rescue [13], and package delivery [17].
[6] provides a comprehensive review of drone-assisted delivery and highlights future research directions. They emphasize that weather conditions, particularly wind, should be accurately accounted for in future studies. Moreover, as noted in [7], existing drone-based research has not fully leveraged advancements from non-drone-related studies, including the use of heterogeneous fleets. These observations underscore a broader challenge: modern facility location problems in UAV applications demand frameworks that can simultaneously address environmental uncertainties and fleet heterogeneity. To bridge this gap, our method explicitly considers a heterogeneous fleet of UAVs and the effect of wind, making it more aligned with real-world applications. This forms another aspect of the motivation for this work in UAV applications.
1.2 Contribution
To address the challenges discussed in the previous section, namely the effects of wind and UAV heterogeneity, we focus on single-facility location problems. These problems have strong theoretical foundations and have been studied in the field of convex analysis, particularly in the context of the Generalized Fermat-Torricelli and Generalized Sylvester problems. To provide a clear foundation for our approach, we first present a brief overview of these classical problems. In the 19th century, English mathematician Sylvester (1814–1897) introduced the smallest enclosing circle problem, which seeks to determine the smallest circle capable of enclosing a given set of points in a plane [41]. This problem is formulated as a minmax location problem and generalized in several ways in the literature. The French mathematician Fermat (1601–1665) proposed an optimization problem that seeks to identify a point in the plane that minimizes the total distance to three specified points. This problem was solved by the Italian mathematician and physicist Torricelli (1608–1647) and is known as the Fermat-Torricelli problem. These problems continue to attract attention both for their mathematical elegance and for their significant practical applications, particularly in location science.
The literature contains several studies on these problems.Variants of the Generalized Sylvester problem have been explored in prior research [23, 31, 29, 30, 34], while the Generalized Fermat-Torricelli problems have been investigated in [27, 24, 28, 21, 19, 15, 33, 12]. Based on the existing literature, we summarize the contributions and generalization aspects of the present work; to the best of our knowledge, these contributions span both theoretical advancements and practical applications, including:
  • Introducing the generalized Sylvester-Fermat-Torricelli (SFT) problem, a novel formulation not previously explored in the literature. The generalized SFT problem is a min–max optimization problem that incorporates multiple Minkowski gauge functions rather than a single one. This generalization provides sufficient flexibility to account for UAV heterogeneity (e.g., varying speeds) and the influence of wind as a source of weather uncertainty. Furthermore, the model extends from individual points to convex sets, which is particularly useful when UAVs are required to reach a target area rather than a specific point. Finally, in this new min–max framework, the maximum of functions—each potentially representing the sum of several Minkowski gauge functions—is minimized. This approach enables modeling of UAV round-trip movements, a feature essential for realistic operational scenarios.
  • Presenting an extended version of the generalized SFT problem, as well as the generalized Fermat–Torricelli and generalized Sylvester problems, incorporating two types of Minkowski gauge functions: the set-based Minkowski gauge function (2.6) and the maximal Minkowski gauge function (3.6). This extension enables modeling of scenarios in which UAVs enter a target area from the point closest to a facility and exit from the point farthest from it. In Section 6, the practical applicability of this extended formulation in real-world operations is demonstrated.
  • Developing a single facility location model for disaster relief using UAVs that explicitly accounts for wind effects and UAV fleet heterogeneity, thereby enhancing its applicability to real-world scenarios.
2 Essential Definitions, Problem Formulation, and Convex Analysis Tools
In this section, we begin by presenting key concepts and tools from convex analysis, which serve as a foundational framework for addressing the challenges posed by the nonsmooth nature of our model introduced in (2.7). These tools will be instrumental in formulating and analyzing the variational properties of the problem, as well as in developing effective solution approaches. Next, we present the essential definitions and the formal problem formulation. We also emphasize the key generalizations introduced in our model, which extend the capabilities beyond those of existing generalized Sylvester problems in the literature. These extensions allow us to address more complex scenarios that are closer to real-world applications.
2.1 Tools of Convex Analysis
In this section, we present key definitions and results from convex analysis that are essential for the subsequent sections.
A subset $\Omega$ of $\mathbb{R}^{q}$ is convex if $\lambda x+(1-\lambda)y\in\Omega$ for all $x,y\in\Omega$ and $\lambda\in(0,1)$. Let $f:\Omega\to\bar{\mathbb{R}}:=(-\infty,\infty]$ be an extended-real-valued function defined on a convex set $\Omega$. Then $f$ is convex on $\Omega$ if
$f\big(\lambda x+(1-\lambda)y\big)\leq\lambda f(x)+(1-\lambda)f(y)\;\mbox{ for all }\;x,y\in\Omega\;\mbox{ and }\;\lambda\in(0,1).$ (2.1)
$f$ is strictly convex if the inequality in (2.1) becomes strict for $x\neq y$.
Let $\Omega\subset\mathbb{R}^{q}$ be a nonempty convex set. For $\bar{x}\in\Omega$, the Normal Cone to $\Omega$ at $\bar{x}$ is
[…]
If $\bar{x}\notin\Omega$, we have $N(\bar{x};\Omega)=\emptyset$ by convention.
The following proposition demonstrates that convexity is preserved under certain important operations (the proof and details are provided in Proposition 1.44 of [22]).
Proposition 2.1.
Let $f_{i}:\mathbb{R}^{q}\to\overline{\mathbb{R}}$ be convex functions for all $i=1,\ldots,m$. Then the following functions are also convex:
  1. 1.
    The scaled function $\alpha f$ for any $\alpha\geq 0$.
  2. 2.
    The sum function $\sum_{i=1}^{m}f_{i}$.
  3. 3.
    The maximum function $\max_{1\leq i\leq m}f_{i}$.
The following lemma shows that the maximum of strictly convex functions is strictly convex.
Lemma 2.2.
Let $f_{i}:\mathbb{R}^{q}\to\bar{\mathbb{R}}$ be strictly convex for all $i=1,2,\dots,m$. Then, the function $f_{m}=\max_{i}f_{i}$ is strictly convex.
Proof.
We begin by proving the result for the case $m=2$, i.e., $f_{m}(x)=\max\{f_{1}(x),f_{2}(x)\}$. For $\in\mathbb{R}^{q}, xyx\neq y$, and $\lambda\in(0,1)$, there are two cases:
Case 1: Suppose $f_{m}(\lambda x+(1-\lambda)y)=f_{1}(\lambda x+(1-\lambda)y)$. Since $f_{1}$ is strictly convex, we have:
$f_{1}(\lambda x+(1-\lambda)y)<\lambda f_{1}(x)+(1-\lambda)f_{1}(y)\leq\lambda f_{m}(x)+(1-\lambda)f_{m}(y),$
so
$f_{m}(\lambda x+(1-\lambda)y)<\lambda f_{m}(x)+(1-\lambda)f_{m}(y).$
Case 2: Similarly, if $f_{m}(\lambda x+(1-\lambda)y)=f_{2}(\lambda x+(1-\lambda)y)$, then the same inequality holds using the strict convexity of $f_{2}$.
Therefore, $f_{m}$ is strictly convex for $m=2$. The general result for $m>2$ follows by induction.
One of the central concepts in convex analysis is the notion of a subgradient, which extends the idea of a derivative to nonsmooth convex functions. Let $f:\mathbb{R}^{q}\to\bar{\mathbb{R}}$ be a convex function, a vector $v\in\mathbb{R}^{q}$ is called a subgradient of $f$ at $\bar{x}$, $\bar{x}\in\operatorname{dom}f$, if the following inequality holds:
$\langle v,x-\bar{x}\rangle\leq f(x)-f(\bar{x})\quad\text{for all }x\in\mathbb{R}^{q}.$ (2.3)
The set of all subgradients of $f$ at $\bar{x}$ is known as the subdifferential of $f$ at $\bar{x}$, and is denoted by $\partial f(\bar{x})$. For completeness, we present two useful rules for computing subdifferential.
Theorem 2.3 (max rule, see Theorem 3.27 in [22]).
Let $f_{i}:\mathbb{R}^{q}\to\overline{\mathbb{R}}$, $i=1,\ldots,m$, be convex functions with $\bar{x}\in\bigcap_{i=1}^{m}\textnormal{dom}\ f_{i}$, and let
$f(x)=\ \max\ \{f_{i}(x)\ |\ i=1,\dots,m\}.$
Then, assuming that $f_{i}$ are continuous at $\bar{x}$ for $i=1,\dots,m$, we have
$\partial f(\bar{x})=\ \textnormal{co }\{\bigcup_{i\in I(\bar{x})}\partial f_{i}(\bar{x})\ \}$
where $I(\bar{x})=\{i=1,\dots,m\ |\ f(\bar{x})=f_{i}(\bar{x})\}$ (the “active” functions at $\bar{x}$).
The theorem says that the subgradient of $f$ at $\bar{x}$ is the convex hull of the subgradients of all functions that achieve the maximum at $\bar{x}$. In the formal statement, dom denotes the domain of a function, i.e., the set of all valid input values for which the function is defined; co denotes the convex hull of a set. Specifically, for a set $\Omega\subset\mathbb{R}^{q}$,
$\operatorname{co}\Omega:=\left\{\sum_{i=1}^{k}\lambda_{i}x_{i}\,\middle|\,x_{i}\in\Omega,\ \lambda_{i}\geq 0,\ \sum_{i=1}^{k}\lambda_{i}=1,\ k\in\mathbb{N}\right\}.$ (2.4)
Theorem 2.4 (The subdifferential of the sum equals the sum of the subdifferential, derived from Theorem 3.18 in [22] by induction.).
Let $f_{i}:\mathbb{R}^{q}\rightarrow\mathbb$$\bar${R}} for $i=1,2,\dots,m$ be convex functions. The following equality holds
$\partial(\sum_{i=1}^{m}f_{i})(\bar{x})=\sum_{i=1}^{m}\partial f_{i}(\bar{x})\ \ \ \textnormal{for every}\ \ \ \bar{x}\in\bigcap_{i=1}^{m}\textnormal{dom}f_{i}$
under the qualification condition:
$\bigcap_{i=1}^{m}\textnormal{ri(dom}f_{i})\neq\emptyset.$
This rule allows us to break down a complicated function into simpler parts in optimization. In the formal statement, ri denotes the relative interior of a set, as defined in (8.2), a subset of points in the set that are in the interior relative to the affine hull of (or space spanned by) the set.
2.2 Problem Formulation
Let $F\subset\mathbb{R}^{n}$ be a nonempty closed convex set. The Minkowski gauge function associated with set $F$, is defined as follows:
$\rho_{F}(x):=\inf\{t\geq 0\ |\ x\in tF\},\ \ x\in\mathbb{R}^{q}.$ (2.5)
The set F is often referred to as the dynamics set. The Minkowski gauge function, which is common in the literature, plays a key role in convex analysis and its applications [22]. However, for our purposes in this paper, we require a more flexible function, as defined below.
Given a nonempty set $\Omega\subset\mathbb{R}^{q}$, we define the set-based Minkowski gauge function associated with the sets $F$ and $\Omega$ as follows:
$\rho_{F}^{\Omega}(x):=\inf\{t\geq 0\ |\ x\in\Omega+tF\},$ (2.6)
where $\Omega+tF:=\{\omega+tf\ |\ \textnormal{for all }\omega\in\Omega\textnormal{ and }f\in F\}$. In the function $\rho_{F}^{\Omega}$, the sets $F$ and $\Omega$ are referred to as the dynamic set and the reference set, respectively. In simple terms, if we view $F$ as a dynamic set representing possible speeds or directions of motion, then $\rho_{F}^{\Omega}(x)$ represents the minimal time required to reach $x$ starting from at least one point $\omega\in\Omega$. Moreover, $\rho_{F}$ can be interpreted in this context by noting that $\rho_{F}(x)=\rho_{F}^{\{0\}}(x)$. The set-based formulation of the Minkowski gauge function has been studied in the literature, where a slightly different notation—denoted by $T_{\Omega}^{F}(x)$ in (8.7)—is commonly used. Intuitively, $T^{F}_{\Omega}(x)$ represents the minimal time required to reach the set $\Omega$ when starting from the point $x$ and moving according to the dynamics defined by the set $F$  [^1] The set of points reachable from the origin in one unit of time.. Since our application is primarily concerned with the minimal time required to reach a point $x$ from the set $\Omega$ under dynamics $F$, we adopt the definition introduced in (2.6) for clarity and consistency. Lemma 8.13 proves that both definitions of the set-based Minkowski gauge function are equivalent.
With these preliminaries in place, we now proceed to formulate the proposed model. Let the index sets $I_{k}\neq\emptyset$ for $k=1,2,\dots,n$ form a partition of the set $I:=\{1,2,\dots,m\}$, such that $\bigcup_{k=1}^{n}I_{k}=I$ and $I_{k}\cap I_{l}=\emptyset$ for all $k\neq l$ with $k,l\in\{1,2,\dots,n\}$. For $i\in I_{k}$ with $k=1,2,\dots n$, let $F_{i}\subset\mathbb{R}^{q}$ be a compact, convex set containing the origin in its interior point (i.e., $0\in\operatorname{int}F_{i}$). Also, let $\Omega_{i}$ and the constraint set $\Omega_{0}$ be nonempty, closed, convex subsets of $\mathbb{R}^{q}$. Then, the generalized Sylvester-Fermat-Torricelli (SFT) problem is proposed:
$\min S(x)\ \ \ s.t.\ \ \ x\in\Omega_{0},$ (2.7)
where
$S(x):=\max\{\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)\ |\ k=1,2,...,n\}.$ (2.8)
In the case where $I_{k}=\{k\}$, the generalized SFT problem reduces to a generalized Sylvester problem, which can be stated as follows:
$\min Y(x)\ \ \ s.t.\ \ \ x\in\Omega_{0},$ (2.9)
where
$Y(x):=\max\{\rho_{F_{i}}^{\Omega_{i}}(x)\ |\ i=1,2,3,\dots,n\}.$ (2.10)
Furthermore, by considering $k=1$ and $|I_{1}|=m$, the generalized SFT problem simplifies to a generalized Fermat–Torricelli problem, which is given by:
$\min T(x)\ \ \ s.t.\ \ \ x\in\Omega_{0},$ (2.11)
where
$T(x):=\sum_{i=1}^{m}\rho_{F_{i}}^{\Omega_{i}}(x).$ (2.12)
In this paper, we primarily focus on the generalized SFT problem, which encompasses both the generalized Sylvester and Fermat–Torricelli problems as special cases. A detailed discussion of these problems is provided separately in the Appendix. It is important to emphasize that, unlike the generalized Sylvester problem in the literature—which typically assumes a single identical dynamic set in the space ($F_{i}=F$ for $i=1,2,\dots,n$)—the generalized SFT problem accommodates multiple dynamic sets. Moreover, another key generalization is the use of $\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)$ instead of a single $\rho_{F_{i}}^{\Omega_{i}}(x)$, which enables modeling of more complex movement patterns, such as round-trip UAV flights under wind conditions between two points.
3 Properties of the Set-Based Minkowski Gauge Function
In this section, we first review several foundational definitions and then develop key properties of the set-based Minkowski gauge function which are:
  1. $\bullet$
    The relationship between the functions $\rho_{F}^{\Omega}$ and $\rho_{F}$, which serves as a foundational tool for deriving further properties and facilitating the computation of $\rho_{F}^{\Omega}$.
  2. $\bullet$
    Convexity of $\rho_{F}^{\Omega}$, which enables the application of tools from convex analysis.
  3. $\bullet$
    Lipschitz continuity of $\rho_{F}^{\Omega}$ which implies that its subdifferential is bounded.
  4. $\bullet$
    Characterization of the subdifferential of $\rho_{F}^{\Omega}$ at both interior points ($x\in\Omega$) and exterior points ($x\notin\Omega$).
  5. $\bullet$
    An upper estimate for the subdifferential of $\rho_{F}^{\Omega}$, offering a useful means for computing its subdifferential at differentiable points.
In the remainder of the section, we introduce the maximal set-based Minkowski gauge function $\bar{\rho}_{F}^{\Omega}$, as defined in (3.6), and establish its fundamental properties.
After presenting the following definitions, we introduce the first property of $\rho_{F}^{\Omega}$. Let $F\subset\mathbb{R}^{q}$ be a nonempty set; the support function associated with $F$ is defined as follows:
$\sigma_{F}(v):=\sup\{\langle v,f\rangle\ |\ f\in F\},\ v\in\mathbb{R}^{q}.$ (3.1)
The $r$-enlargement of a set $\Omega\subset\mathbb{R}^{q}$ is defined as:
$\Omega_{r}:=\{x\in\mathbb{R}^{q}\mid\rho_{F}^{\Omega}(x)\leq r\},\ \ \text{for any }\ r>0.$ (3.2)
As the first property of $\rho_{F}^{\Omega}$, we establish its relationship with the Minkowski gauge function.
Proposition 3.1 (The relationship between $\rho_{F}^{\Omega}$ and $\rho_{F}$).
Let $F\subset\mathbb{R}^{q}$ be a nonempty closed convex set. Then, for a nonempty set $\Omega\subset\mathbb{R}^{q}$, we have:
$\rho_{F}^{\Omega}(x)=\inf_{\omega\in\Omega}\rho_{F}(x-\omega).$ (3.3)
Proof.
First, consider the case where $\rho_{F}^{\Omega}(x)<\infty$. Let $\epsilon>0$, and suppose $\rho_{F}^{\Omega}(x)=p$. Choose $t$ such that $p\leq t<p+\epsilon$ and $x\in\Omega+tF$. Then, there exists $\omega\in\Omega$ such that $x-\omega\in tF$, which implies $\rho_{F}(x-\omega)\leq t<p+\epsilon$. Since $\rho_{F}^{\Omega}(x)=p$, it follows that
$\inf_{\omega\in\Omega}\rho_{F}(x-\omega)\leq\rho_{F}(x-\omega)<\rho_{F}^{\Omega}(x)+\epsilon.$
Taking the limit as $\epsilon\to 0$, we obtain
$\inf_{\omega\in\Omega}\rho_{F}(x-\omega)\leq\rho_{F}^{\Omega}(x).$
For the reverse inequality, suppose $\inf_{\omega\in\Omega}\rho_{F}(x-\omega)=q$. Let $\alpha>0$, and choose $d$ and $\omega$ such that $q\leq d<q+\alpha$ as well as $x-\omega\in dF$. Then, $x\in\omega+dF\subset\Omega+dF$. Hence, $\rho_{F}^{\Omega}(x)\leq d<q+\alpha$. Taking the limit as $\alpha\to 0$, we get
$\rho_{F}^{\Omega}(x)\leq\inf_{\omega\in\Omega}\rho_{F}(x-\omega).$
Combining both inequalities, we conclude that
$\rho_{F}^{\Omega}(x)=\inf_{\omega\in\Omega}\rho_{F}(x-\omega).$
Finally, consider the case where $\rho_{F}^{\Omega}(x)=\infty$. Then, the set $\{t\geq 0\mid x\in\Omega+tF\}=\emptyset$. This implies that for every $\omega\in\Omega$, the set $\{t\geq 0\mid x-\omega\in tF\}=\emptyset$, so $\rho_{F}(x-\omega)=\infty$. Therefore,
$\inf_{\omega\in\Omega}\rho_{F}(x-\omega)=\infty=\rho_{F}^{\Omega}(x).$
For the next property, we prove the convexity of the set-based Minkowski gauge function under the convexity of the reference set in the following proposition.
Proposition 3.2 (Convexity of $\rho_{F}^{\Omega}$).
Let $F\subset\mathbb{R}^{q}$ be a nonempty closed convex set. If $\Omega\subset\mathbb{R}^{q}$ is nonempty and convex, then $\rho_{F}^{\Omega}(.)$ in (2.6) is a convex function.
Proof.
Given $\epsilon>0$ and select $t_{1}$ and $t_{2}$ such that $\rho_{F}^{\Omega}(x_{i})\leq t_{i}<\rho_{F}^{\Omega}(x_{i})+\epsilon$ as well as $x_{i}\in\Omega+t_{i}f_{i}$ where $f_{i}\in F$ for $i=1,2$. Since $x_{i}-t_{i}f_{i}\in\Omega$ (for $i=1,2$) and convexity of $\Omega$, we have $\lambda(x_{1}-t_{1}f_{1})+(1-\lambda)(x_{2}-t_{2}f_{2})\in\Omega$, for $\lambda\in(0,1)$.
Hence,
$\lambda x_{1}+(1-\lambda)x_{2}\in\Omega+\lambda t_{1}f_{1}+(1-\lambda)t_{2}f_{2}\subset\Omega+\lambda t_{1}F+(1-\lambda)t_{2}F.$
By convexity of $F$, $\lambda t_{1}F+(1-\lambda)t_{2}F=(\lambda t_{1}+(1-\lambda)t_{2})F$, so we have
$\lambda x_{1}+(1-\lambda)x_{2}\in\Omega+(\lambda t_{1}+(1-\lambda)t_{2})F\ \Rightarrow\ \rho_{F}^{\Omega}(\lambda x_{1}+(1-\lambda)x_{2})\leq\lambda t_{1}+(1-\lambda)t_{2}.$
By arbitrary choice of $\epsilon$, we have
$\lambda t_{1}+(1-\lambda)t_{2}<\lambda\rho_{F}^{\Omega}(x_{1})+(1-\lambda)\rho_{F}^{\Omega}(x_{2})+\epsilon\Rightarrow\rho_{F}^{\Omega}(\lambda x_{1}+(1-\lambda)x_{2})\leq\lambda\rho_{F}^{\Omega}(x_{1})+(1-\lambda)\rho_{F}^{\Omega}(x_{2}).$
The following proposition establishes the Lipschitz continuity of the function $\rho_{F}^{\Omega}$—a key property for our analysis.
Proposition 3.3 (Lipschitz Continuity of $\rho_{F}^{\Omega}$).
Let $F\subset\mathbb{R}^{q}$ be a nonempty, closed, convex set with $0\in$ int $F$. Then for a nonempty $\Omega\subset\mathbb{R}^{q}$, $\rho_{F}^{\Omega}(\cdot)$ is Lipschitz continuous.
Proof.
The function $\rho_{F}$ is subadditive and positively homogeneous (Theorem 6.14 in [22] provides its proof). Therefore, for any $x,y\in\mathbb{R}^{q}$, we have:
$\rho_{F}(x-\omega)\leq\rho_{F}(x-y)+\rho_{F}(y-\omega)\quad\text{for all }\omega\in\Omega,$
$\Rightarrow\inf_{\omega\in\Omega}\rho_{F}(x-\omega)\leq\rho_{F}(x-y)+\inf_{\omega\in\Omega}\rho_{F}(y-\omega)$
$\Rightarrow\inf_{\omega\in\Omega}\rho_{F}(x-\omega)-\inf_{\omega\in\Omega}\rho_{F}(y-\omega)\leq\rho_{F}(x-y)$
$\Rightarrow\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(y)\leq\rho_{F}(x-y).$
Also, for the function $\rho_{F}$, there exists a constant $l\geq 0$ such that $\rho_{F}(x)\leq l\|x\|$ for all $x\in\mathbb{R}^{q}$, as shown in Proposition 6.18 of [22]. Hence, we have:
$\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(y)\leq\rho_{F}(x-y)\leq l\|x-y\|.$
After establishing Lipschitz continuity, we proceed to characterize the subdifferential of the function $\rho_{F}^{\Omega}(x)$ at points $x\in\Omega$ (interior points).
Proposition 3.4 (Subdifferential of $\rho_{F}^{\Omega}$ at interior points).
Let $F\subset\mathbb{R}^{q}$ be a nonempty compact convex set with $0\in\textnormal{int}F$, and let $\Omega$ be a nonempty convex set in $\mathbb{R}^{q}$. Then, the subdifferential of $\rho_{F}^{\Omega}(\cdot)$ at $\bar{x}\in\Omega$ is given by
$\partial\rho_{F}^{\Omega}(\bar{x})=N(\bar{x};\Omega)\cap\mathcal{V}$
where $\mathcal{V}$ is defined as
$\mathcal{V}:=\{v\in\mathbb{R}^{q}\mid\sigma_{F}(v)\leq 1\}.$
Proof.
Fix any $v\in\partial\rho_{F}^{\Omega}(\bar{x})$ and for $x\in\Omega$, $\rho_{F}^{\Omega}(x)=0$, we get the inequlities,
$\langle v,x-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})=0.$
Therefore, $v\in N(\bar{x};\Omega)$. For $x\not\in\Omega$, there exists $t>0$ and $f\in F$ such that $\bar{x}+tf=x$ which means $\rho_{F}^{\Omega}(x)\leq t$. Then we have:
$\langle v,\bar{x}+tf-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-0\leq t.$
Therefore, $\sigma_{F}(v)\leq 1$, and $\partial\rho_{F}^{\Omega}(\bar{x})\subseteq N(\bar{x};\Omega)\cap\mathcal{V}$.
To prove the inverse inclusion, fix any $v\in N(\bar{x};\Omega)\cap\mathcal{V}$. Given $\epsilon>0$, for $x\in\mathbb{R}^{q}$, select $t>0$ such that $\rho_{F}^{\Omega}(x)\leq t<\rho_{F}^{\Omega}(x)+\epsilon$. Then, find $\omega\in\Omega$ and $f\in F$ such that $\omega+tf=x$. Therefore,
$\langle v,\omega+tf-\bar{x}\rangle=\langle v,\omega-\bar{x}\rangle+t\langle v,f\rangle\leq t<\rho_{F}^{\Omega}(x)+\epsilon=\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})+\epsilon,$
and $\langle v,x-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})$ as $\epsilon\rightarrow 0$. Hence, $v\in\partial\rho_{F}^{\Omega}(\bar{x})$, and $N(\bar{x};\Omega)\cap\mathcal{V}\subseteq\partial\rho_{F}^{\Omega}(\bar{x}).$
We are now prepared to present the second theorem regarding the subdifferential of the $\rho_{F}^{\Omega}(x)$ function at points $x\notin\Omega$ (exterior points).
Proposition 3.5 (Subdifferential of $\rho_{F}^{\Omega}$ at exterior points).
Let $F\subset\mathbb{R}^{q}$ be a nonempty compact convex set with $0\in\textnormal{int}F$, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty convex set. Suppose $\rho_{F}^{\Omega}(\bar{x})=r>0$, where $\bar{x}\notin\Omega$ and $\bar{x}\in\operatorname{dom}\rho_{F}^{\Omega}$. Then the subdifferential of $\rho_{F}^{\Omega}$ at $\bar{x}$ is given by:
$\partial\rho_{F}^{\Omega}(\bar{x})=N(\bar{x};\Omega_{r})\cap\Upsilon\ \ \textnormal{with}\ \ \Upsilon:=\{v\in\mathbb{R}^{q}\ |\ \sigma_{F}(v)=1\}$
Proof.
Fix $v\in\partial\rho_{F}^{\Omega}(\bar{x})$. We consider three cases:
First, suppose $x\notin\Omega_{r}$. Then, there exist $t>0$ and $f\in F$ such that $\bar{x}+tf=x$, which implies $\rho_{F}^{\Omega}(x)\leq t+r$ by Lemma 8.7. Hence, we obtain the inequality:
$\langle v,\bar{x}+tf-\bar{x}\rangle=\langle v,tf\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})\leq t+r-r=t,$
which implies $\langle v,f\rangle\leq 1$, and thus $\sigma_{F}(v)\leq 1$.
Second, suppose $x\in\Omega_{r}$. Then the following inequality holds:
$\langle v,x-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})\leq 0,$
which implies $v\in N(\bar{x},\Omega_{r})$.
Third, consider the case $x\in\Omega$. For any $\epsilon>0$, there exist $x\in\Omega$ and $f\in F$ such that $x+df=\bar{x}$, where $r\leq d<r+\epsilon$. Then, we have:
$\langle v,x-\bar{x}\rangle=\langle v,\bar{x}-df-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})$
$\Rightarrow r\leq\langle v,df\rangle<(r+\epsilon)\langle v,f\rangle$
$\Rightarrow\frac{r}{r+\epsilon}<\langle v,f\rangle\leq\sigma_{F}(v).$
As $\epsilon\to 0$, we obtain $\sigma_{F}(v)\geq 1$. Hence, we conclude:
$v\in N(\bar{x};\Omega_{r})\cap\left\{v\in\mathbb{R}^{q}\mid\sigma_{F}(v)\leq 1\right\}\cap\left\{v\in\mathbb{R}^{q}\mid\sigma_{F}(v)\geq 1\right\}=N(\bar{x};\Omega_{r})\cap\Upsilon,$
which implies $\partial\rho_{F}^{\Omega}(\bar{x})\subseteq N(\bar{x};\Omega_{r})\cap\Upsilon$.
For the opposite inclusion, take any $v\in N(\bar{x};\Omega_{r})\cap\Upsilon$, we need to show that the following inequality holds:
$\langle v,x-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})\ \ \textnormal{for all}\ \ x\in\mathbb{R}^{q}.$ (3.4)
By applying Proposition 3.4, we know that $v\in\partial\rho_{F}^{\Omega_{r}}(\bar{x})$, which gives:
$\langle v,x-\bar{x}\rangle\leq\rho_{F}^{\Omega_{r}}(x)-\rho_{F}^{\Omega_{r}}(\bar{x}).$
If $x\notin\Omega_{r}$, we can use Lemma 8.6, and then:
$\langle v,x-\bar{x}\rangle\leq\rho_{F}^{\Omega_{r}}(x)=\rho_{F}^{\Omega}(x)-r=\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x}),$
which implies $v\in\partial\rho_{F}^{\Omega}(\bar{x})$.
In the other case, suppose $x\in\Omega_{r}$ or $\rho_{F}^{\Omega}(x)=p\leq r$. Since $\sigma_{F}(v)=1$, for any $\epsilon>0$, we can find $f\in F$ such that $\langle v,f\rangle\geq 1-\epsilon$. Using Lemma 8.7, we obtain $y\in\Omega_{r}$, where y=x+(rp)fy=x+(r-p)f. Furthermore, since $v\in N(\bar{x};\Omega_{r})$, it follows that $\langle v,y-\bar{x}\rangle\leq 0$. Therefore:
$\langle v,x-(p-r)f-\bar{x}\rangle=\langle v,x-\bar{x}\rangle-(p-r)\langle v,f\rangle\leq 0$
$\Rightarrow\langle v,x-\bar{x}\rangle\leq(p-r)\langle v,f\rangle\leq(1-\epsilon)(p-r)=(1-\epsilon)(\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})).$
Hence, $v\in\partial\rho_{F}^{\Omega}(\bar{x})$ as $\epsilon\rightarrow 0$, and we obtain $N(\bar{x};\Omega_{r})\cap\Upsilon\subseteq\partial\rho_{F}^{\Omega}(\bar{x})$. ∎
Now, we need an important definition for the following proposition. Let $F\subset\mathbb{R}^{q}$ be a nonempty compact convex set with $0\in\textnormal{int}F$, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty closed convex set. Then, the generalized projection from a point $x\in\mathbb{R}^{q}$ to the set $\Omega$, denoted by $\Pi_{F}(x;\Omega)$, is defined as follows:
$\Pi_{F}(x;\Omega):=\{\omega\in\Omega\ |\ \rho_{F}^{\Omega}(x)=\rho_{F}(x-w)\}.$ (3.5)
The nonemptiness of $\Pi_{F}(x;\Omega)$ is established in Proposition 8.1.
The following proposition provides an upper estimate for the subdifferential of the function $\rho_{F}^{\Omega}$, which is useful for computing the subdifferential at points where the function is differentiable.
Proposition 3.6 (Subdifferential estimation for $\rho_{F}^{\Omega}$).
Let $F\subset\mathbb{R}^{q}$ be a nonempty compact convex set with $0\in\textnormal{int }F$, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty closed convex set. The subdifferential of the set-based Minkowski gauge function at $\bar{x}\in\textnormal{dom}\rho_{F}^{\Omega}$ can be estimated as follows:
$\partial\rho_{F}^{\Omega}(\bar{x})\subseteq\partial\rho_{F}(\bar{x}-\bar{\omega}),$
where $\bar{\omega}\in\Pi_{F}(\bar{x};\Omega)$. If, in addition, $\bar{x}\in\textnormal{int(dom}\rho_{F}^{\Omega})$ and $\rho_{F}$ is differentiable at $\bar{x}-\bar{\omega}$, then the inclusion becomes an equality:
$\partial\rho_{F}^{\Omega}(\bar{x})=\partial\rho_{F}(\bar{x}-\bar{\omega})=\{\nabla\rho_{F}(\bar{x}-\bar{\omega})\}.$
Proof.
Let $v\in\partial\rho_{F}^{\Omega}(\bar{x})$. Then, by Proposition 3.1, it follows that:
$\langle v,x-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})=\inf_{\omega\in\Omega}\rho_{F}(x-\omega)-\inf_{\omega\in\Omega}\rho_{F}(\bar{x}-\omega)\leq\rho_{F}(x-\bar{\omega})-\rho_{F}(\bar{x}-\bar{\omega}).$
Therefore $v\in\partial\rho_{F}(\bar{x}-\bar{\omega})$ and $\partial\rho_{F}^{\Omega}(\bar{x})\subseteq\partial\rho_{F}(\bar{x}-\bar{\omega})$.
For the final part of the proposition, we know that $\rho_{F}^{\Omega}$ is convex due to Proposition 3.2. Therefore, by Theorem 23.4 in [37], its subdifferential is nonempty at $\bar{x}\in\operatorname{int}(\operatorname{dom}\rho_{F}^{\Omega})$. Furthermore, since $\rho_{F}$ is differentiable at $\bar{x}-\bar{\omega}$, the subdifferential $\partial\rho_{F}(\bar{x}-\bar{\omega})$ reduces to a singleton (see Theorem 5.2 in [22]). It then follows that
$\partial\rho_{F}^{\Omega}(\bar{x})=\partial\rho_{F}(\bar{x}-\bar{\omega}).$
Finally, by Theorem 25.1 in [37], we have $\partial\rho_{F}(\bar{x}-\bar{\omega})=\{\nabla\rho_{F}(\bar{x}-\bar{\omega})\}.$
The next proposition establishes the relationship between $\rho_{F}^{\Omega}$ and its positive scalar multiples, demonstrating how positive coefficients and weights can be incorporated into the dynamic set.
Proposition 3.7 (Relationship between $\rho_{F}^{\Omega}$ and its positive scalar multiple).
Let $F\subset\mathbb{R}^{q}$ be a nonempty closed convex set, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty set. Then, for $\lambda>0$ we have:
$\lambda\rho^{\Omega}_{F}(x)=\rho^{\Omega}_{\frac{F}{\lambda}}(x).$
Proof.
First, consider the case where $\rho_{F}^{\Omega}(x)<\infty$. Given $\epsilon>0$, choose $t$ such that $\rho_{F}^{\Omega}(x)\leq t<\rho_{F}^{\Omega}(x)+\epsilon$. Then, for any $\epsilon$, there exist $\omega\in\Omega$ and $f\in F$ such that $x=\omega+tf$. We can rewrite this as $x=\omega+\lambda t\cdot\frac{f}{\lambda},$ which implies that $x\in\Omega+(\lambda t)\frac{F}{\lambda}$. Therefore, $\rho_{\frac{F}{\lambda}}^{\Omega}(x)\leq\lambda t<\lambda(\rho_{F}^{\Omega}(x)+\epsilon).$ Letting $\epsilon\to 0$ yields $\rho_{\frac{F}{\lambda}}^{\Omega}(x)\leq\lambda\rho_{F}^{\Omega}(x).$ For the reverse inequality, let $\alpha>0$, and choose $d$ such that $\rho_{\frac{F}{\lambda}}^{\Omega}(x)\leq d<\rho_{\frac{F}{\lambda}}^{\Omega}(x)+\alpha$. Then, for any $\alpha$, there exist $\omega\in\Omega$ and $\frac{f}{\lambda}\in\frac{F}{\lambda}$ such that $x=\omega+d\cdot\frac{f}{\lambda}.$ This implies $x\in\Omega+\frac{d}{\lambda}F$, and therefore, $\rho_{F}^{\Omega}(x)\leq\frac{d}{\lambda}<(\frac{1}{\lambda})(\rho_{\frac{F}{\lambda}}^{\Omega}(x)+\alpha)$. Multiplying through by $\lambda$ gives $\lambda\rho_{F}^{\Omega}(x)<\rho_{\frac{F}{\lambda}}^{\Omega}(x)+\alpha$. Letting $\alpha\to 0$ yields $\lambda\rho_{F}^{\Omega}(x)\leq\rho_{\frac{F}{\lambda}}^{\Omega}(x)$. Combining both inequalities, we conclude that $\rho_{\frac{F}{\lambda}}^{\Omega}(x)=\lambda\rho_{F}^{\Omega}(x)$.
Next, consider the case where $\rho_{F}^{\Omega}(x)=\infty$, therefore, $\{t\geq 0\ |\ x\in\Omega+tF\}=\emptyset$. This implies that the set $\{t\geq 0\ |\ x\in\Omega+t\frac{F}{\lambda}\}=\emptyset$, so
$\lambda\rho_{F}^{\Omega}(x)=\infty=\rho_{\frac{F}{\lambda}}^{\Omega}(x).$
In the final part of this section, we define a new function that will be useful in practical applications.
Let $F\subset\mathbb{R}^{n}$ be a nonempty closed bounded convex set, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty bounded set; then, for a point $x\in\mathbb{R}^{q}$, the Maximal Set-based Minkowski Gauge (MSMG) function is defined as:
$\bar{\rho}_{F}^{\Omega}(x):=\inf\{t\geq 0\ |\ x-\Omega\subseteq tF\}.$ (3.6)
In simple terms, the set $x-\Omega$ includes all vectors that start from a point in $\Omega$ and end at $x$. If we view $F$ as a dynamic set, then $\bar{\rho}_{F}^{\Omega}(x)$ represents the minimal time required to reach $x$ from the farthest point in $\Omega$. This function has been investigated in the literature, where a slightly different notation, $C_{\Omega}^{F}(x)$, as defined in (8.8), is commonly employed. Lemma 8.14 formalizes the connection between $C_{\Omega}^{F}(x)$ and the MSMG function. The application of the MSMG function will be presented in Section 6. The following proposition demonstrates the relationship between the MSMG function and $\rho_{F}$.
Proposition 3.8 (Relationship between $\bar{\rho}_{F}^{\Omega}$ and $\rho_{F}$).
Let $F\subset\mathbb{R}^{q}$ be a nonempty closed bounded convex set with $0\in\textnormal{int }F$. Then, for a nonempty bounded set $\Omega\subset\mathbb{R}^{q}$, we have:
$\bar{\rho}_{F}^{\Omega}(x)=\sup_{\omega\in\Omega}\rho_{F}(x-\omega).$ (3.7)
Proof.
First, consider the case where $\bar{\rho}_{F}^{\Omega}(x)=0$. This implies that $\Omega=\{x\}$. Consequently, $\sup_{\omega\in\Omega}\rho_{F}(x-\omega)=0$, and the equality in (3.7) holds.
Next, consider the case $\bar{\rho}_{F}^{\Omega}(x)>0$, and let $\epsilon>0$ be arbitrary. Set $t=\bar{\rho}_{F}^{\Omega}(x)-\epsilon$, so that $t>0$. By the definition of $\bar{\rho}_{F}^{\Omega}(x)$, there exists $\omega\in\Omega$ such that $x-\omega\notin tF$. This implies that $t<\rho_{F}(x-\omega)\leq\sup_{\omega\in\Omega}\rho_{F}(x-\omega)$. Since $t\to\bar{\rho}_{F}^{\Omega}(x)$ as $\epsilon\to 0$, it follows that $\bar{\rho}_{F}^{\Omega}(x)\leq\sup_{\omega\in\Omega}\rho_{F}(x-\omega)$.
For the reverse inequality, let $d$ be such that $\bar{\rho}_{F}^{\Omega}(x)\leq d<\bar{\rho}_{F}^{\Omega}(x)+\alpha$, where $\alpha>0$ is arbitrary. By the definition of $\bar{\rho}_{F}^{\Omega}(x)$, this implies that $x-\omega\in dF$ for all $\omega\in\Omega$. Hence $\rho_{F}(x-\omega)\leq d$ for all $\omega\in\Omega$. Therefore $\sup_{\omega\in\Omega}\rho_{F}(x-\omega)\leq d<\bar{\rho}_{F}^{\Omega}(x)+\alpha$. As $\alpha\to 0$, $\sup_{\omega\in\Omega}\rho_{F}(x-\omega)\leq\bar{\rho}_{F}^{\Omega}(x)$.
The following theorem outlines key properties of the MSMG function that are fundamental to its practical implementation.
Theorem 3.9 (Properties of the MSMG function).
Let $F\subset\mathbb{R}^{q}$ be a nonempty compact convex set with $0\in\textnormal{int}F$, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty compact convex set. For the MSMG function, we have:
  • i)
    The MSMG function is convex.
  • ii)
    The MSMG function is Lipschitz continuous.
  • iii)
    The subdifferential of the MSMG function can be estimated at $\bar{x}\in\textnormal{dom}\bar{\rho}_{F}^{\Omega}$ as follows:
    $\partial\rho_{F}(\bar{x}-\omega^{\prime})\subseteq\partial\bar{\rho}_{F}^{\Omega}(\bar{x}),$
    where $\omega^{\prime}\in\bar{\Pi}_{F}(\bar{x};\Omega):=\{\omega\in\mathbb{R}\ |\ \bar{\rho}_{F}^{\Omega}(\bar{x})=\rho_{F}(\bar{x}-\omega)\}$.
  • iv)
    The subdifferential of the MSMG function can be obtained at $\bar{x}\in\textnormal{dom }\bar{\rho}_{F}^{\Omega}$ as follows:
    $\partial\bar{\rho}_{F}^{\Omega}(\bar{x})=\textnormal{co}\left\{\bigcup_{\omega^{\prime}\in\bar{\Pi}_{F}(\bar{x};\Omega)}N(\omega^{\prime};\bar{x}-rF)\cap\Upsilon^{\prime}\right\},$
    where $\Upsilon^{\prime}=\{v\in\mathbb{R}^{q}\ |\ \sigma_{-F}(v)=1\}$, $\omega^{\prime}\in\bar{\Pi}_{F}(\bar{x};\Omega)$ as defined in $(iii)$, and $r=\rho_{-F}^{\{\bar{x}\}}(\omega^{\prime})$.
Proof.
For $(i)$, using Proposition 3.2, we obtain that $\rho_{F}$ is convex. By applying Propositions 3.8 and 2.1, we conclude that $\bar{\rho}_{F}^{\Omega}(\cdot)$ is convex. For $(ii)$, it follows from the subadditivity of the function $\rho_{F}$ (Theorem 6.14 in [22] provides its proof) we have:
$\rho_{F}(x-\omega)\leq\rho_{F}(x-y)+\rho_{F}(y-\omega)\quad\text{for all }\omega\in\Omega,$
$\Rightarrow\sup_{\omega\in\Omega}\rho_{F}(x-\omega)\leq\rho_{F}(x-y)+\sup_{\omega\in\Omega}\rho_{F}(y-\omega)$
$\Rightarrow\sup_{\omega\in\Omega}\rho_{F}(x-\omega)-\sup_{\omega\in\Omega}\rho_{F}(y-\omega)\leq\rho_{F}(x-y)$
$\Rightarrow\bar{\rho}_{F}^{\Omega}(x)-\bar{\rho}_{F}^{\Omega}(y)\leq\rho_{F}(x-y).$
Also, for the function $\rho_{F}$, there exists a constant $l\geq 0$ such that $\rho_{F}(x)\leq l\|x\|$, as shown in Proposition 6.18 of [22]. Then, we have:
$\bar{\rho}_{F}^{\Omega}(x)-\bar{\rho}_{F}^{\Omega}(y)\leq\rho_{F}(x-y)\leq l(x-y).$
Therefore, $\bar{\rho}_{F}^{\Omega}(\cdot)$ is Lipschitz continuous due to the arbitrary choice of $x$ and $y$.
For (iiiiii), given the continuity of $\rho_{F}$ and the compactness of $\Omega$, we obtain $\bar{\Pi}_{F}(\bar{x};\Omega)\neq\emptyset$ by applying the Weierstrass Theorem. Take any $v\in\partial\rho_{F}(\bar{x}-\omega^{\prime})$; then we have:
$\langle v,x-\bar{x}\rangle\leq\rho_{F}(x-\omega^{\prime})-\rho_{F}(\bar{x}-\omega^{\prime})\leq\sup_{\omega\in\Omega}\rho_{F}(x-\omega)-\sup_{\omega\in\Omega}\rho_{F}(\bar{x}-\omega)=\bar{\rho}_{F}^{\Omega}(x)-\bar{\rho}_{F}^{\Omega}(\bar{x}).$
Therefore, $v\in\partial\bar{\rho}_{F}^{\Omega}(\bar{x})$, which implies that $\partial\rho_{F}(\bar{x}-\omega^{\prime})\subseteq\partial\bar{\rho}_{F}^{\Omega}(\bar{x}).$
To prove statement $(iv)$, we first show that
$\sup_{\omega\in\Omega}\rho_{-F}^{\{\bar{x}\}}(\omega)=\rho_{-F}^{\{\bar{x}\}}(\omega^{\prime}),$
where $\omega^{\prime}:=\bar{\Pi}_{F}(\bar{x};\Omega)$. To this end, we apply Proposition 3.1 for $\rho_{-F}^{\{\bar{x}\}}(\omega)$, which gives $\sup_{\omega\in\Omega}\rho_{-F}^{\{\bar{x}\}}(\omega)=\sup_{\omega\in\Omega}\rho_{-F}(\omega-\bar{x}).$ Using Lemma 8.12, we have $\rho_{-F}(\omega-\bar{x})=\rho_{F}(\bar{x}-\omega),$. By the definition of $\bar{\Pi}_{F}(\bar{x};\Omega)$ and Proposition 3.8, we know that $\omega^{\prime}\in\bar{\Pi}_{F}(\bar{x};\Omega)$ satisfies
$\rho_{F}(\bar{x}-\omega^{\prime})=\sup_{\omega\in\Omega}\rho_{F}(\bar{x}-\omega).$
Equivalently,
$\rho_{-F}(\omega^{\prime}-\bar{x})=\sup_{\omega\in\Omega}\rho_{-F}(\omega-\bar{x}),$
and hence
$\sup_{\omega\in\Omega}\rho_{-F}^{\{\bar{x}\}}(\omega)=\rho_{-F}^{\{\bar{x}\}}(\omega^{\prime}).$ (3.8)
Using Proposition 8.10 and the equality in (3.8), we have:
$\bar{\rho}_{F}^{\Omega}(\bar{x})=\sup_{\omega\in\Omega}\rho_{-F}^{\{\bar{x}\}}(\omega)=\rho_{-F}^{\{\bar{x}\}}(\omega^{\prime}).$
Then, by Proposition 3.5, we can write:
$\partial\rho_{-F}^{\{\bar{x}\}}(\omega^{\prime})=N(\omega^{\prime};\bar{x}-rF)\cap\Upsilon^{\prime},$
Finally, by applying Proposition 2.3, we obtain:
$\partial\bar{\rho}_{F}^{\Omega}(\bar{x})=\textnormal{co}\left\{\bigcup_{\omega^{\prime}\in\bar{\Pi}_{F}(\bar{x};\Omega)}N(\omega^{\prime};\bar{x}-rF)\cap\Upsilon^{\prime}\right\}.$
4 Existence and Uniqueness of Optimal Solutions in the Generalized SFT Problem and Its Weighted Variant
At the beginning of this section, we define a preliminary term. Given two points $x,y\in\mathbb{R}^{q}$, the line passing through them is defined by
$L(x,y):=\{\lambda x+(1-\lambda)y\mid\lambda\in\mathbb{R}\}.$ (4.1)
In the following Theorem, we provide the condition under which the existance of the solution will be proved.
Theorem 4.1 (Existence of solutions to the generalized SFT problem).
In the setting of the generalized SFT problem (2.7), an optimal solution exists if at least one of the sets $\Omega_{0}$ or $\Omega_{i}$ (for some $i\in I_{k}$, $k=1,\dots,n$) is bounded.
Proof.
The sublevel set for $\lambda\geq 0$ and the function $S(x)$ defined in (2.8) is given by
$\mathcal{L}_{\lambda}(S):=\{x\in\Omega_{0}\mid S(x)\leq\lambda\},$
which implies that $\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)\leq\lambda$ for all $k=1,2,\dots,n$, whenever $x\in\Omega_{0}$. Therefore, for the set
$A:=\left\{x\in\Omega_{0}\mid\rho_{F_{i}}^{\Omega_{i}}(x)\leq\lambda\ \text{for all}\ i\in I_{k},\ k=1,2,\dots,n\right\},$
we have $\mathcal{L}_{\lambda}(S)\subseteq A$. For a closed set $\Omega_{i}$, the condition $\rho_{F_{i}}^{\Omega_{i}}(x)\leq\lambda$ implies that $x\in\Omega_{i}+\lambda F_{i}$. Therefore, by the definition of $A$, it follows that
$A=\Omega_{0}\cap\left(\bigcap_{k=1}^{n}\left(\bigcap_{i\in I_{k}}(\Omega_{i}+\lambda F_{i})\right)\right).$
Hence, if $\Omega_{i}$ is bounded for some $i\in I_{k}$ with $k=1,2,\dots,n$, then the set $A$ is bounded. Since $A$ is bounded, $\mathcal{L}_{\lambda}(S)$ is also bounded. Then, by Corollary 7.10 in [22], the existence of an optimal solution is guaranteed.
In the following theorem, under certain assumptions, we establish the uniqueness of the optimal solution in the generalized SFT problem.
Theorem 4.2 (Uniqueness of solutions to the generalized SFT problem).
In the setting of the generalized SFT problem (2.7), a unique solution exists under the following conditions:
  1. i)
    The sets $F_{i}$ and $\Omega_{i}$ are strictly convex for $i\in I_{k}$ with $k=1,2,\dots,n$.
  2. ii)
    For each index set $I_{k}$ with $k\in\{1,2,\dots,n\}$, among the sets $\Omega_{i}$ with $i\in I_{k}$, there exists $\Omega_{j}$ such that for all distinct points $x,y\in\Omega_{0}$, the line $L(x,y)$ does not intersect $\Omega_{j}$; that is, $L(x,y)\cap\Omega_{j}=\emptyset$.
  3. iii)
    At least one of the set $\Omega_{0}$ or the sets $\Omega_{i}$ for $i\in I_{k}$ with $k=1,\dots,n$ is bounded.
Proof.
Based on $(iii)$ and Theorem 4.1, we know that there is an optimal solution. Based on $(i)$, $(ii)$, and Proposition 8.18, each $\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)$ is strictly convex for $k=1,2,\dots,n$. It follows from Lemma 2.2 that $S(x)$ is strictly convex, thereby completing the proof.
Motivated by prior works that studied the weighted generalized Fermat–Torricelli and Sylvester problems [25, 35, 36], we conclude this section by presenting and discussing the weighted formulation of the generalized SFT problem, given by (considering the setting of the problem in (2.7)):
$\min\ \tilde{S}(x)\quad\text{subject to}\quad x\in\Omega_{0},$ (4.2)
where, for weights $\omega_{i}>0$ for all $i\in I_{k}$ and $k=1,2,\dots,n$, the objective function is defined as:
$\tilde{S}(x):=\max\{\sum_{i\in I_{k}}\omega_{i}\rho_{F_{i}}^{\Omega_{i}}(x)\ |\ k=1,2,\dots,n\}.$ (4.3)
It is worth noting that the original formulation of the generalized SFT problem in (2.7) inherently includes its weighted counterpart. By applying Proposition 3.7, we obtain
$\omega_{i}\rho_{F_{i}}^{\Omega_{i}}(x)=\rho_{\frac{F_{i}}{\omega_{i}}}^{\Omega_{i}}(x),$
which allows us to express the weighted version in the same form as the original problem:
$\min\ S(x)\quad\text{subject to}\quad x\in\Omega_{0},$ (4.4)
where
$S(x):=\max\{\sum_{i\in I_{k}}\rho_{\tilde{F}_{i}}^{\Omega_{i}}(x)\ |\ k=1,2,\dots,n\},$ (4.5)
and $\tilde{F}_{i}:=F_{i}/\omega_{i}$.
5 An Extended Formulation of the Generalized SFT Problem
In this section, we give an extended version of the generalized SFT problem that incorporates both the set-based Minkowski gauge and the MSMG functions. In addition to the setting of the generalized SFT problem (2.7) consider the following settings. Let $\bar{m}\in\mathbb{N}$ and the index sets $J_{k}\neq\emptyset$ for $k=1,2,\dots,n$ form a partition of the set $J:=\{1,2,\dots,\bar{m}\}$, such that $\bigcup_{k=1}^{n}J_{k}=J$ and $J_{k}\cap J_{l}=\emptyset$ for all $k\neq l$ with $k,l\in\{1,2,\dots,n\}$. For $j\in J_{k}$ with $k=1,2,\dots n$, let $\bar{F}_{j}\subset\mathbb{R}^{q}$ be a compact, convex set containing the origin in its interior point (i.e., $0\in\operatorname{int}\bar{F}_{j}$). Let $\bar{F}_{j}\subset\mathbb{R}^{q}$ be a compact, convex set containing the origin in its interior point (i.e., $0\in\operatorname{int}\bar{F}_{j}$). Also, let $\Theta_{j}$ be nonempty compact convex subsets of $\mathbb{R}^{q}$. Then, the following problem is formulated as an extension of the generalized SFT problem:
$\min\hat{S}(x)\ \ \ s.t.\ \ \ x\in\Omega_{0},$ (5.1)
where
$\hat{S}(x):=\max\{\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)+\sum_{j\in J_{k}}\bar{\rho}_{\bar{F}_{j}}^{\Theta_{j}}(x)\ |\ k=1,2,...,n\}.$ (5.2)
The application of the extended generalized SFT problem (5.1) is demonstrated in a real-world scenario in Section 6.
In the case $k=1$, $|I_{1}|=m$, and $|J_{1}|=\bar{m}$ in problem (5.1), the extended version of the generalized Fermat-Torricelli problem is given by:
$\min\hat{T}(x)\ \ \ s.t.\ \ \ x\in\Omega_{0},$ (5.3)
where
$\hat{T}(x):=\sum_{i=1}^{m}\rho_{F_{i}}^{\Omega_{i}}(x)+\sum_{j=1}^{\bar{m}}\bar{\rho}_{\bar{F}_{j}}^{\Theta_{j}}(x).$ (5.4)
For the generalized Fermat-Torricelli problem in (2.11), a similar problem was studied by [12]. However, the problem proposed in (5.3) represents a new extension of the Fermat-Torricelli problem, and its application in disaster relief operations using UAVs is discussed in Section 6.3.
Finally, in the setting of the problem (5.1), we introduce the extended formulation of the generalized Sylvester problem:
$\min\hat{Y}(x)\quad\text{s.t.}\quad x\in\Omega_{0},$ (5.5)
where
$\hat{Y}(x)=\{\mathcal{Y}_{1}(x),\mathcal{Y}_{2}(x)\},$ (5.6)
with
$\mathcal{Y}_{1}(x)=\max_{i=1,2,\dots,m}\rho_{F_{i}}^{\Omega_{i}}(x),\quad\mathcal{Y}_{2}(x)=\max_{j=1,2,\dots,\bar{m}}\bar{\rho}_{F_{j}}^{\Theta_{j}}(x).$ (5.7)
It should be noted that both problems proposed in (2.9) and (5.5) constitute novel generalizations of the Sylvester problem. Their application in disaster relief operations using UAVs is discussed in Sections 6.2 and 6.3.
The convexity of the functions $\hat{S}(x)$, $\hat{T}(x)$, and $\hat{Y}(x)$—defined in (5.2), (5.4), and (5.6), respectively—follows directly from Proposition 3.2, Proposition 2.1, and Theorem 3.9. Furthermore, their subdifferentials and subgradients can be derived using Propositions 3.4, 3.5, 3.6, together with Theorems 3.9, 2.3, and 2.4.
6 Application of the Generalized SFT Problem to Optimizing Facility Location Decisions in Disaster Relief Operations
After an earthquake, some people may be trapped under debris but still alive. If they have access to their cell phones, one of the most effective ways they can help themselves is by sending their location to family members or posting it on social media. This can greatly increase their chances of being found and rescued. However, earthquakes often damage or destroy communication systems, such as cell towers and internet cables. As a result, even if someone has a working phone, it may not be able to send messages or connect to the internet.
In such cases, UAVs can play an important role. By flying over the damaged area, UAVs can provide wireless connections or DTNs, allowing trapped individuals to send their location information (see Fig. 1). This support can make rescue operations faster and more effective, and help save more lives. Despite their advantages, UAVs also present several challenges. They are limited by short battery life and are vulnerable to environmental conditions, especially high winds. Moreover, rescue teams often deploy a heterogeneous fleet, consisting of UAVs with varying speeds, ranges, and operational capabilities, which adds complexity to coordination and mission planning.
In this section, we examine several scenarios in which a truck serves as a mobile station for deploying and recharging UAVs. The primary objective is to determine the optimal location of the truck (see Fig. 1) to minimize the transition time—the period when the UAV is traveling between the service area and the mobile station, resulting in a temporary loss of network service. This problem can be formulated as a single-facility location problem, which has been extensively studied in the literature using the $\ell_{1}$ and $\ell_{2}$ norms. However, to model a more realistic scenario, we incorporate the effect of wind, which prevents us from using these common norms directly. Using the $\ell_{2}$ norm, for example, would ignore the wind’s influence, leading to a simplification that can result in significant time loss. To minimize transition time, we use the set-based Minkowski gauge function (2.6) that captures the effects of wind, which are inherent in real-world conditions.
In the first scenario, a rescue team utilizes a single UAV, and we show how the generalized Fermat-Torricelli problem can be applied to minimize transition time. In the second scenario, the rescue team operates a heterogeneous fleet of UAVs, and we illustrate how the generalized Sylvester problem can effectively address the complexities that arise. Finally, we consider the generalized SFT problem in a scenario where the heterogeneous fleet of UAVs is first deployed to the affected area and then returns to the truck for recharging and data transfer. This setup highlights the operational challenges and the importance of strategic truck placement in dynamic rescue missions.
Figure 1: Providing a wireless network or DTNs after an earthquake is one of the important applications of UAVs in disaster relief
6.1 Application of the Generalized Fermat-Torricelli Poblem
Owing to constraints such as limited budget, it is often the case that a rescue team operates with only a single UAV. In this section, we examine an operation in which the rescue team employs a UAV to provide DTN support for three predetermined areas. First, the UAV is dispatched from the mobile station (e.g., a truck) and covers the first area. After collecting data, the UAV returns to the station to recharge and transfer the collected data to the rescue team. This pattern is then repeated for the remaining two areas. To determine the optimal location for the truck—minimizing transition time under windy conditions—we apply the generalized Fermat-Torricelli problem, a variation of the generalized SFT problem. We present a step-by-step solution to illustrate how the generalized SFT problem is applied in this context.
To formulate the stated challenge (corresponding to Case 1 in Table 1) as a generalized Fermat-Torricelli problem (2.11), we begin by specifying the reference and dynamic sets. The reference sets are:
$\Omega_{0}$ $=\mathbb{R}^{2},$
$\Omega_{1}$ $=\left\{(x,y)\in\mathbb{R}^{2}\mid\max\{|x-30|,|y-350|\}\leq 15\right\},$
$\Omega_{2}$ $=\left\{(x,y)\in\mathbb{R}^{2}\mid\max\{|x-210|,|y-10|\}\leq 15\right\},$
$\Omega_{3}$ $=\left\{(x,y)\in\mathbb{R}^{2}\mid\max\{|x-550|,|y-200|\}\leq 15\right\}.$
The dynamic sets in the absence of wind, corresponding to the UAV’s nominal speed $r$ (m/s), are represented by the set:
$F_{0}=\left\{(f_{1},f_{2})\in\mathbb{R}^{2}\mid f_{1}^{2}+f_{2}^{2}\leq r^{2}\right\}\ \ \textnormal{for}\ \ i=1,2,3.$
Taking wind into account, with wind vector $\mathcal{S}\in\mathbb{R}^{2}$ and $\mathcal{S}=(s_{1},s_{2})$, the effective dynamic set—representing the UAV’s velocity under wind conditions—is given by
$F_{i}=F=\left\{(f_{1},f_{2})\in\mathbb{R}^{2}\mid(f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq r^{2}\right\}\ \ \textnormal{for}\ \ i=1,2,3.$ (6.1)
Figure 2: Case 1 in Table 1: (a) Approximated optimal location of the truck when the wind effect is neglected. (b) Dynamic set with $\|\mathcal{S}\|=0$. (c) Approximated optimal location of the truck under windy conditions. (d) Dynamic sets of the UAV under windy conditions with $\mathcal{S}=(-0.6,0.6)$. The black arrow indicates the wind direction
a)Refer to caption     b)Refer to caption c)Refer to caption     d)Refer to caption
Dispatching the UAV from the mobile station involves movement in the opposite direction, so the corresponding dynamic set is given by $F_{i}$ for $i=1,2,3$, as stated in Proposition 8.9. We assume that the wind vector satisfies the condition $\left\|\mathcal{S}\right\|_{2}<r$, ensuring that the UAV can still operate effectively under the given windy condition. Therefore, the generalized Fermat-Torricelli problem is represented as (for simplicity, we assume $\bar{F}=-F$):
$\min_{x\in\Omega_{0}}T(x)\ \ \textnormal{where}\ \ T(x)=\sum_{i=1}^{3}\rho_{F_{i}}^{\Omega_{i}}(x)+\sum_{i=1}^{3}\rho_{-F_{i}}^{\Omega_{i}}(x)=\sum_{i=1}^{3}\rho_{F}^{\Omega_{i}}(x)+\sum_{i=1}^{3}\rho_{\bar{F}}^{\Omega_{i}}(x),$ (6.2)
where $x=(x_{1},x_{2})\in\mathbb{R}^{2}$. To solve the problem (6.2), we divide the solution approach into several steps.
Step 1: Obtaining the Minkowski gauge function $\rho_{F}(x):$
We begin by computing the Minkowski gauge function (2.5). The function $\rho_{F}(x)$ is obtained as the optimal value of the following optimization problem:
$\begin{split}&\min\ t\\ \text{s.t.}\quad&(f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}=r^{2},\\ &x_{1}=tf_{1},\\ &x_{2}=tf_{2},\\ &t\geq 0.\end{split}$ (6.3)
According to Proposition 8.8, the first constraint in (6.3) enforces that the velocity vector $(f_{1},f_{2})$ lies on the boundary of the dynamic set $F$. To derive a closed-form expression for $\rho_{F}(x)$, we can substitute $f_{1}=x_{1}/t$ and $f_{2}=x_{2}/t$ into the first constraint in (6.3), yielding:
$\begin{split}&\ (\frac{x_{1}}{t}-s_{1})^{2}+(\frac{x_{2}}{t}-s_{2})^{2}=r^{2}\\ \Rightarrow&\ (x_{1}-ts_{1})^{2}+(x_{2}-ts_{2})^{2}=r^{2}t^{2}\\ \Rightarrow&\ t^{2}(s_{1}^{2}+s_{2}^{2}-r^{2})-2(s_{1}x_{1}+s_{2}x_{2})t+(x_{1}^{2}+x_{2}^{2})=0\end{split}$ (6.4)
By solving the quadratic equation obtained from (6.4), we have:
$\rho_{F}(x)=t=\frac{(s_{1}x_{1}+s_{2}x_{2})-\sqrt{x_{1}^{2}(r^{2}-s_{2}^{2})+x_{2}^{2}(r^{2}-s_{1}^{2})+2s_{1}s_{2}x_{1}x_{2}}}{s_{1}^{2}+s_{2}^{2}-r^{2}}$ (6.5)
The value of $t$ obtained from (6.5) is the optimal solution to the problem in (6.3), and is obviously nonnegative.
Step 2: Obtaining the subdiferential of $\rho_{F}(x)$ and $\rho_{F}^{\Omega}(x)$ functions:
In this step, we compute the subdifferentials of the functions $\rho_{F}(x)$ and $\rho_{F}^{\Omega}(x)$, which are necessary for implementing the subgradient algorithm (Algorithm 1)—a well-known method in convex optimization—in order to find the minimizer of the problem in (6.2). Let $\alpha=\frac{1}{s_{1}^{2}+s_{2}^{2}-r^{2}}$, and define
$\beta(x_{1},x_{2})=x_{1}^{2}(r^{2}-s_{2}^{2})+x_{2}^{2}(r^{2}-s_{1}^{2})+2s_{1}s_{2}x_{1}x_{2}.$
Then, for $x\neq 0$, the gradient of the function $\rho_{F}(x)$ in (6.5) is given by:
$\nabla\rho_{F}(x)=\alpha(s_{1}-\frac{x_{1}(r^{2}-s_{2}^{2})+s_{1}s_{2}x_{2}}{\sqrt{\beta(x_{1},x_{2})}},s_{2}-\frac{x_{2}(r^{2}-s_{1}^{2})+s_{1}s_{2}x_{1}}{\sqrt{\beta(x_{1},x_{2})}}).$ (6.6)
Therefore, the subdifferential of the function $\rho_{F}^{\Omega_{i}}$ can be expressed as follows:
$\partial\rho_{F}^{\Omega_{i}}(x)=\begin{cases}N(x;\Omega_{i})\cap\mathcal{V}&\text{if }x\in\Omega_{i},\\ \{\nabla\rho_{F}(x-\omega_{i})\}&\text{if }x\notin\Omega_{i},\end{cases}\quad\text{for }i=1,2,3,$ (6.7)
where $\omega_{i}\in\Pi_{F}(x;\Omega_{i})$, and $\mathcal{V}$ is defined in Proposition 3.4. Note that in (6.7), for $x\notin\Omega_{i}$, $x-\omega_{i}\neq 0$ and the subdifferential is obtained using Proposition 3.6. For $x\in\Omega_{i}$, we have $x=\Pi_{F}(x;\Omega_{i})$, and (6.6) does not provide the gradient at the origin. Therefore, we apply Proposition 3.4 to compute the subdifferential of $\rho_{F}^{\Omega_{i}}(x)$ at $x\in\Omega_{i}$.
Table 1: Experimental Results for Generalized Fermat-Torricelli Cases
[…]
No. (Table 4) $(\mathcal{S}=(s_{1},s_{2}))$ $x^{\ast}$ $Z_{N}(s)$ $x^{\ast}$ $Z_{I}(s)$ Imp(s) Imp($\%$)
1 Info 1 (0.6,0.6)(-0.6,0.6) (252,117) 4039 (242,52) 3959 80 1.98
2 Info 1 $(0.6,0.6)$ (252,117) 3626 (225,171) 3575 53 1.46
3 Info 1 $(0.7,0.3)$ (252,117) 2746 (228,108) $27$ 19 0.00
4 Info 1 (0.1,0.9)(-0.1,0.9) (252,117) 5214 (350,162) 5037 177 3.39
5 Info 1 (0.4,0.1)(0.4,-0.1) $(252,117)$ 1605 $(247,102)$ 1604 1 0.00
6 Info 2 (0.8,0.3)(0.8,-0.3) (604,265) 8305 (581,101) 8148 157 1.89
7 Info 2 (0,0.2)(0,-0.2) (604,265) 2728 (605,270) 2728 0 0.00
8 Info 3 (0.4,0.8)(0.4,-0.8) $(166,48)$ 3103 (175,35) 3061 42 1.35
9 Info 3 (0.3,0) (166,48) 896 (164,44) 896 0 0.00
10 Info 3 $(0.4,0.8)$ (166,48) 2742 (176,84) 2597 145 5.28
$Z_{N}$ and $Z_{I}$ denote the objective function values (transition time in seconds) for the cases ignoring and considering wind, respectively. Absolute Imp ($Z_{N}-Z_{I}$) and Relative Imp ($(Z_{N}-Z_{I})/Z_{N}$) denote the absolute and relative improvements.
Step 3: Obtaining the optimal solution by subgradient algorithm:
To determine the optimal location of the mobile station, we apply the subgradient algorithm (Algorithm 1) to find the minimizer of the generalized Fermat–Torricelli problem (6.2). According to the structure of the subgradient algorithm, we need to compute a subgradient of $T(x)$. To do so, we calculate the subdifferential of $T(x)$ by applying the subdifferential sum rule (Theorem 2.4), and we obtain:
$\partial T(x)=\sum_{i=1}^{3}\partial\rho_{F}^{\Omega_{i}}(x)+\sum_{i=1}^{3}\partial\rho_{\bar{F}}^{\Omega_{i}}(x).$ (6.8)
Therefore, at each iteration $k$ of the subgradient algorithm, a subgradient $v^{k}\in\partial T(x^{k})$ is computed as
$v^{k}=\sum_{i=1}^{3}v_{i}^{k}+\sum_{i=1}^{3}\bar{v}_{i}^{k},$ (6.9)
where each $v_{i}^{k}\in\partial\rho_{F}^{\Omega_{i}}(x^{k})$ and $\bar{v}_{i}^{k}\in\partial\rho_{\bar{F}}^{\Omega_{i}}(x^{k})$ is determined based on the formula in (6.7) (taking into account that $0\in N(x^{k};\Omega_{i})\cap\mathcal{V}$):
$v_{i}^{k}=\begin{cases}0&\text{if }x^{k}\in\Omega_{i},\\ \nabla\rho_{F}(x^{k}-\omega_{i}^{k})&\text{if }x^{k}\notin\Omega_{i},\end{cases}\quad\text{for }i=1,2,3,$ (6.10)
with $\omega_{i}^{k}\in\Pi_{F}(x^{k};\Omega_{i})$. Also, we have:
$\bar{v}_{i}^{k}=\begin{cases}0&\text{if }x^{k}\in\Omega_{i},\\ \nabla\rho_{\bar{F}}(x^{k}-\bar{\omega}_{i}^{k})&\text{if }x^{k}\notin\Omega_{i},\end{cases}\quad\text{for }i=1,2,3,$ (6.11)
where $\bar{\omega}_{i}^{k}\in\Pi_{\bar{F}}(x^{k};\Omega_{i})$. Note that by solving the problem in (6.3) with $F$ replaced by $\bar{F}$, we can compute both $\rho_{\bar{F}}(x)$ and its gradient $\nabla\rho_{\bar{F}}(x)$.
To illustrate the procedure, we focus on Case 1 in Table 1 and evaluate $v^{k}\in\partial T(x^{k})$ for $k=1$, where the dynamic sets are specified by Info 1 in Table 4 and the wind vector is given by $\mathcal{S}=(-0.6,0.6)$. Therefore, the dynamic sets for this UAV, corresponding to back-and-forth movements, are given by:
$\begin{split}F=\{(f_{1},f_{2})\in\mathbb{R}^{2}|(f_{1}+0.6)^{2}+(f_{2}-0.6)^{2}\leq 1\},\\ \bar{F}=\{(f_{1},f_{2})\in\mathbb{R}^{2}|(f_{1}-0.6)^{2}+(f_{2}+0.6)^{2}\leq 1\}.\end{split}$ (6.12)
Let the initial point be $x^{1}=(100,100)$, and consider the computation of $v_{1}^{1}$. Since $x^{1}\notin\Omega_{1}$, the formula in (6.10) implies that it is sufficient to compute $\nabla\rho_{F}(x^{1}-\omega_{1}^{1})$. With $\omega_{1}^{1}=(45,335)$, we have
$x^{1}-\omega_{1}^{1}=(100-45,\,100-335)=(55,-235).$
Applying the formula in (6.6), we obtain
$\alpha=-3.57,\quad\beta(55,-235)=46586,\quad\text{and}\quad\nabla\rho_{F}(55,-235)=$
$(-3.57)\Bigg(-0.6-\frac{55(0.64)+(-235)(-0.36)}{\sqrt{46586}},\;0.6-\frac{(-235)(0.64)+55(-0.36)}{\sqrt{46586}}\Bigg)=(4.12,-4.95).$
Using the same procedure, we obtain:
[…]
Finally, $v^{1}$ can be computed using (6.9), yielding
$v^{1}=(-6.66,1.43).$
Following this procedure, we compute $v^{k}$ at each iteration. By choosing $\alpha^{k}=\tfrac{1}{k}$ and applying Algorithm 1, we obtain an approximate solution to the generalized Fermat-Torricelli problem (6.2).
For Case 1 in Table 1, we first simplify the problem by neglecting the wind effect and considering the dynamic sets for the back-and-forth movement as $F=\bar{F}=F_{0}$, as illustrated in Fig. 2(b). In this scenario, the approximate solution is $(252,117)$, shown in Fig. 2(a). If this solution is applied in a real-world scenario with wind, the resulting transition time is 4039 seconds. When the wind effect is incorporated, the dynamic sets are defined as in (6.12), illustrated in Fig. 2(d). The optimal solution then shifts to $(242,52)$, depicted in Fig. 2(c), reducing the objective value to 3960 seconds. It should be noted that, since the sets $\Omega_{i}$ for $i=1,2,3$ are not strictly convex, the uniqueness of the problem’s solution cannot be guaranteed, as stated in Proposition 8.18.
Table 1 presents ten cases with varying settings and parameters to evaluate the efficiency of the generalized Fermat–Torricelli problem in accounting for wind effects. For simplicity, all cases are conducted on a small scale, with the UAV’s nominal speed set to $1\text{(m/s)}$; these parameters can be appropriately scaled to reflect the real scenarios. The results indicate that, for single-UAV operations, the generalized Fermat–Torricelli problem reduces the transition time by approximately $3.3\%$.
Figure 3: Case 2 in Table 2: (a) Approximated optimal location of the truck when the wind effect is neglected. (b) Dynamic sets with $\|\mathcal{S}\|=0$. (c) Approximated optimal location of the truck under windy conditions. (d) Dynamic sets of the UAVs under windy conditions with $\mathcal{S}=(-0.6,0.6)$. The black arrow indicates the wind direction
a)Refer to caption     b)Refer to caption c)Refer to caption     d)Refer to caption
Algorithm 1 subgradient Algorithm
1:Input: Initial point $x^{1}\in\Omega_{0}$, step size sequence $\{\alpha^{k}>0\}$, maximum iterations $K\in\mathbb{N}$.
2:Initialize: Set $k=1$ ($k\in\mathbb{N}$).
3:while $k<K$ do
4:  Choose a subgradient $v^{k}\in\partial f(x^{k})$.
5:  Update: $x^{k+1}=\Pi(x^{k}-\alpha^{k}v^{k};\Omega_{0})$ $\triangleright$ $\Pi(.;\Omega_{0})$ denotes Euclidean projection onto $\Omega_{0}$
6:  $k\leftarrow k+1$
7:end while
8:Output: $x^{K}$
Note. $\{\alpha^{k}\}_{k=1}^{\infty}$ should satisfies (1) $\alpha^{k}\to 0$ and (2) $\sum_{k=1}^{\infty}\alpha^{k}=\infty$
6.2 Application of the Generalized Sylvester Problem
In the context of Section 6.1, we consider a scenario in which the rescue team deploys a heterogeneous fleet of three UAVs. These UAVs are launched at the earliest possible time following a disaster, and the objective is to determine the optimal location for a mobile station (truck) to collect them as quickly as possible. Timely retrieval of the UAVs enables rapid access to critical data, which is essential for effective response and decision-making in emergency situations. We consider this problem by the following setting (corresponding to Case 2 in Table 2):
$\Omega_{0}$ $=\mathbb{R}^{2},$
$\Omega_{1}$ $=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-30)^{2}}{10^{2}}+\frac{(y-350)^{2}}{10^{2}}\leq 1\},$
$\Omega_{2}$ $=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-210)^{2}}{10^{2}}+\frac{(y-10)^{2}}{10^{2}}\leq 1\},$
$\Omega_{3}$ $=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-550)^{2}}{10^{2}}+\frac{(y-200)^{2}}{10^{2}}\leq 1\}.$
For dynamic sets, we have:
$\ F_{1}$ $=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}+0.6)^{2}+(f_{2}-0.6)^{2}\leq 2^{2}\},$
$\ F_{2}$ $=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}+0.6)^{2}+(f_{2}-0.6)^{2}\leq 1^{2}\},$
$\ F_{3}$ $=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}+0.6)^{2}+(f_{2}-0.6)^{2}\leq 3^{2}\}.$
We set 1(m/s) as the nominal speed of the slowest UAV (which is represented by $F_{2}$); consequently, the rescue team is equipped with two additional UAVs that are twice and three times faster, respectively. In addition, we consider the wind vector as $\mathcal{S}=(-0.6,0.6)$.
To determine the optimal location of the mobile station, we formulate the problem as the following generalized Sylvester problem, which is a variant of the generalized SFT problem:
$\min_{x\in\Omega_{0}}Y(x)\ \ \ \textnormal{where}\ \ \ Y(x):=\max\{\rho_{F_{i}}^{\Omega_{i}}(x)\ |\ \ i=1,2,3\}.$ (6.13)
To apply Algorithm 1 for approximating a minimizer of the problem in (6.13), we need to compute a subgradient of $Y(x)$ at each iteration $k$. This is done by following the procedure described in Section 6.1 to compute $v_{i}^{k}\in\partial\rho_{F_{i}}^{\Omega_{i}}(x^{k})$. Subsequently, $v^{k}\in\partial Y(x^{k})$ is obtained using the formula derived from Theorem 2.3:
$\partial Y(x^{k})=\operatorname{co}\left\{\bigcup_{i\in I(x^{k})}\partial\rho_{F_{i}}^{\Omega_{i}}(x^{k})\right\}.$
Moreover, the uniqueness of the solution is guaranteed by Theorem 8.17.
Table 2: Experimental Results for Generalized Sylvester Cases
[…]
No. (Table 4) $(\mathcal{S}=(s_{1},s_{2}))$ $x^{\ast}$ $Z_{N}(s)$ $x^{\ast}$ $Z_{I}(s)$ Imp(s) Imp($\%$)
1 Info 4 (0.7,0.7)(-0.7,0.7) (175,138) 242 (95,228) 124 100 41.3
2 Info 4 (0.6,0.6)(-0.6,0.6) (175,138) 212 (108,216) 124 88 41.5
3 Info 4 (0.5,0.5)(-0.5,0.5) (175,138) 190 (116,202) 124 66 34.7
4 Info 4 (0.4,0.4)(-0.4,0.4) (175,138) 171 (122,188) 124 47 27.4
5 Info 4 (0.3,0.3)(-0.3,0.3) (175,138) 156 (146,179) 124 32 20.5
6 Info 4 (0.2,0.2)(-0.2,0.2) (175,138) 143 (151,163) 123 20 13.9
7 Info 4 (0.1,0.1)(-0.1,0.1) (175,138) 133 (175,139) 123 10 7.51
8 Info 4 (0.6,0.6) (175,138) 163 (261,217) 125 38 23.3
9 Info 4 (0.6,0.6)(0.6,-0.6) (175,138) 721 (265,64) 126 595 82.5
10 Info 4 (0.6,0.6)(-0.6,-0.6) (175,138) 481 (118,67) 128 353 73.3
11 Info 5 (0.6,0.6)(0.6,-0.6) (239,163) 1401 (371,40) 212 1189 84.8
12 Info 6 (0.6,0.6)(0.6,-0.6) (331,227) 525 (386,140) 315 210 40.0
13 Info 6 (0.8,0)(-0.8,0) (331,227) 1156 (161,124) 309 847 73.2
14 Info 6 $(0.8,0)$ (331,227) 432 (474,311) 306 126 29.1
$Z_{N}$ and $Z_{I}$ denote the objective function values (transition time in seconds) for the cases ignoring and considering wind, respectively. Absolute Imp ($Z_{N}-Z_{I}$) and Relative Imp ($(Z_{N}-Z_{I})/Z_{N}$) denote the absolute and relative improvements.
Initially, we consider the case where the wind effect is ignored ($\mathcal{S}=0$) and the dynamic sets are illustrated in Fig. 3(b). The approximated solution in this case is $(175,138)$, as shown in Fig. 3(a), which corresponds to a transition time of $212$ seconds in the real scenario under windy conditions. When the wind effect is taken into account with $\mathcal{S}=(-0.6,0.6)$, the dynamic sets are modified as depicted in Fig. 3(d), yielding the approximated optimal location $(108,216)$, as illustrated in Fig. 3(c). The transition time for $(108,216)$ is $124$ seconds, representing a $41.5\%$ reduction in transition time. This demonstrates the effectiveness of the generalized Sylvester problem in addressing this challenge.
To evaluate the performance of the generalized Sylvester problem in mitigating wind effects, 14 cases with varying settings and parameters are considered, as summarized in Table 2. For simplicity, all cases are performed on a small scale, assigning 11(m/s) to the slowest UAV’s nominal speed; these values can be scaled to represent real-world scenarios. The results demonstrate that the generalized Sylvester problem significantly reduces transition time, achieving an approximate $84\%$ improvement. In Cases 1 to 7 of Table 2, all parameters are identical except for the length of the wind vector. The results of these cases are illustrated in Fig. 5(b), which highlights the impact of wind speed. As expected, at higher wind speeds, applying the generalized Sylvester problem resulted in greater efficiency and a larger reduction in transition time.
6.3 Application of the Generalized SFT Problem
In the setting described in Section 6.2, the rescue team first determines the optimal location for the mobile station and then deploys a fleet of three UAVs. After completing their missions, the UAVs return to the mobile station for data transfer and recharging. To address this challenge while accounting for wind conditions and fleet heterogeneity (corresponding to Case 1 in Table 3), we employ the following generalized SFT model:
$\min_{x\in\Omega_{0}}S(x),$ (6.14)
where
$S(x):=\max\{\rho_{F_{i}}^{\Omega_{i}}(x)+\rho_{\bar{F}_{i}}^{\Omega_{i}}(x)\ |\ i=1,2,3\},$ (6.15)
such that, $\bar{F}_{i}=-F_{i}$.
To compute $\partial S(x)$, we apply Theorems 2.3 and 2.4, yielding:
$\partial S(x)=\operatorname{co}\left\{\bigcup_{i\in I(x)}\left(\partial\rho_{F_{i}}^{\Omega_{i}}(x)+\partial\rho_{\bar{F}_{i}}^{\Omega_{i}}(x)\right)\,\middle|\,i=1,2,3\right\},$ (6.16)
Figure 4: Case 1 in Table 3: (a) Approximated optimal location of the truck when the wind effect is neglected. (b) Dynamic sets with $\|\mathcal{S}\|=0$. (c) Approximated optimal location of the truck under windy conditions. (d) Dynamic sets of the UAVs under windy conditions with $\mathcal{S}=(-0.6,0.6)$. The black arrow indicates the wind direction
a)Refer to caption     b)Refer to caption c)Refer to caption     d)Refer to caption
where $\partial\rho_{F_{i}}^{\Omega_{i}}(x)$ and $\partial\rho_{\bar{F}_{i}}^{\Omega_{i}}(x)$ is obtained by procedure outlined in Section 6.1. The subgradient algorithm (Algorithm 1) can be utilized to compute an optimal solution to problem (6.14). In continuation, we consider a more challenging case of this scenario, which has a real-world application. In this case, UAVs are deployed from a point $x$ (the location of the mobile station), first travel to the nearest point in $\Omega_{i}$, and then return from the farthest point in $\Omega_{i}$ back to $x$. Such cases commonly arise when UAVs are used to cover regions $\Omega_{i}$ for tasks such as data collection or aerial photography of affected areas, employing coverage path planning algorithms (see, for example, [13, 5, 4, 16]). Then, in the worst-case situation, UAVs complete their coverage at the farthest point and then must return to the mobile station for recharging and data transfer. This case can be formulated through the extended version of the generalized SFT problem (5.1), as follows:
$\min_{x\in\Omega_{0}}\hat{S}(x),$ (6.17)
where
$\hat{S}(x):=\max\{\rho_{\bar{F}_{i}}^{\Omega_{i}}(x)+\bar{\rho}_{F_{i}}^{\Omega_{i}}(x)\ |\ i=1,2,3\}$ (6.18)
such that $\bar{F}_{i}=-F_{i}$.
To obtain $\partial S(x)$, we apply Theorems 2.3 and 2.4, yielding:
$\partial\hat{S}(x)=\operatorname{co}\left\{\bigcup_{i\in I(x)}\left(\partial\rho_{\bar{F}_{i}}^{\Omega_{i}}(x)+\partial\bar{\rho}_{F_{i}}^{\Omega_{i}}(x)\right)\,\middle|\,i=1,2,3\right\}.$ (6.19)
We use Algorithm 1 to find an approximate minimizer of this problem. At each step of the algorithm, a subgradient from $\partial\hat{S}(x)$ is required. To obtain this, we compute $\partial\rho_{\bar{F}_{i}}^{\Omega_{i}}(x)$ by following the procedure described in Section 6.1. Moreover, for simplicity, instead of fully calculating $\partial\bar{\rho}_{F_{i}}^{\Omega_{i}}(x)$, we employ its estimation provided in Theorem 3.9. In particular, we have
$\nabla\rho_{F_{i}}(x-\omega^{\prime}_{i})\subseteq\partial\bar{\rho}_{F_{i}}^{\Omega_{i}}(x),$
where $\omega_{i}^{\prime}\in\bar{\Pi}_{F_{i}}(x;\Omega_{i})$, as defined in Theorem 3.9. Finally, by (6.19), obtaining a subgradient of $\partial\hat{S}(x)$ becomes straightforward.
For the problem in (6.17), associated with Case 1 in Table 3, the first step is to identify the approximate optimal location of the truck under the condition $\|\mathcal{S}\|=0$, with the corresponding dynamic sets shown in Fig. 4(b). The obtained solution is $(171,134)$, as shown in Fig. 4(a), with a corresponding transition time of $905\,\text{s}$ under real conditions with wind. When incorporating the wind vector $\mathcal{S}=(-0.6,0.6)$, the dynamic sets change accordingly, as illustrated in Fig. 4(d). In this case, the approximate optimal location of the truck shifts to $(189,56)$, as shown in Fig. 4(c), with a reduced transition time of $401\,\text{s}$. This represents a $55.6\%$ improvement, demonstrating the effectiveness of the generalized SFT problem (extended variant) in addressing weather uncertainty. These results highlight that neglecting wind effects can result in a time loss of nearly $500\,\text{s}$—an inefficiency that may be critical in time-sensitive rescue operations.
Table 3: Experimental Results for Generalized SFT Cases (Extended Version)
[…]
No. (Table 4) $(\mathcal{S}=(s_{1},s_{2}))$ $x^{\ast}$ $Z_{N}(s)$ $x^{\ast}$ $Z_{I}(s)$ Imp(s) Imp($\%$)
1 Info 4 (0.6,0.6)(-0.6,0.6) (171,134) 905 (189,56) 401 504 55.6
2 Info 4 $(0.6,0.6)$ (171,134) 637 (175,78) 342 295 46.3
3 Info 4 (0.6,0.6)(-0.6,-0.6) (171,134) 581 $(181,85)$ 340 241 41.4
4 Info 4 (0.6,0.6)(0.6,-0.6) (171,134) 798 (214,88) 394 404 50.6
5 Info 4 $(0.85,0.0)$ (171,134) 528 (183,91) 344 184 34.8
6 Info 4 $(0.0,0.85)$ (171,134) 961 (160,50) 389 572 59.5
7 Info 4 (0.85,0.0)(-0.85,0.0) (171,134) 565 (188,88) 348 217 38.4
8 Info 4 (0.0,0.85)(0.0,-0.85) (171,134) 844 (156,62) 384 460 54.5
9 Info 5 (0.4,0.8) (351,234) 1842 (300,90) 610 1232 66.8
10 Info 5 $(0.1,0.1)$ (351,234) 450 $(333,223)$ 441 9 2.0
11 Info 5 (0.1,0.1)(-0.1,-0.1) (351,234) 447 (347,231) 444 3 0.0
12 Info 5 (0.1,0.1)(0.1,-0.1) (351,234) 448 $(336,225)$ 443 5 1.11
13 Info 5 (0.1,0.1)(-0.1,0.1) (351,234) 453 (340,225) 444 9 1.98
14 Info 6 $(0.8,0)$ (333,220) 1442 (234,138) 829 613 42.5
15 Info 6 $(0.56,0.56)$ (333,220) 1648 (225,124) 877 771 46.7
16 Info 6 $(0,0.8)$ (333,220) 1364 (246,136) 819 545 39.9
17 Info 6 (0.56,0.56)(-0.56,0.56) (333,220) 981 (269,178) 734 247 25.1
18 Info 6 (0.8,0)(-0.8,0) (333,220) 1375 (245,141) 822 553 40.2
19 Info 6 (0.56,0.56)(-0.56,-0.56) (333,220) 1564 (235,135) 869 695 44.4
20 Info 6 (0,0.8)(0,-0.8) (333,220) 1307 (235,153) 819 488 37.3
21 Info 6 (0.56,0.56)(0.56,-0.56) (333,220) 987 (267,178) 734 253 25.6
22 Info 7 $(0.69,0.4)$ (2265,1459) 4148 (2805,1335) 2028 2120 51.1
23 Info 7 $(0.56,0.56)$ (2265,1459) 3840 (2878,1295) 2064 1776 46.2
24 Info 7 $(0.8,0.0)$ (2265,1459) 4117 (2673,1456) 1955 2162 52.5
$Z_{N}$ and $Z_{I}$ denote the objective function values (transition time in seconds) for the cases ignoring and considering wind, respectively. Absolute Imp ($Z_{N}-Z_{I}$) and Relative Imp ($(Z_{N}-Z_{I})/Z_{N}$) denote the absolute and relative improvements.
If, in the case modeled by the problem in (6.17), UAVs are deployed at the earliest time and the objective is to determine the optimal location of the mobile station for collecting them, as described in Section 6.2, the extended generalized Sylvester problem (5.5) can be used to optimally locate the mobile station. Furthermore, if the rescue team employs a single UAV instead of three (as in the scenario presented in Section 6.1), the UAV travels from the mobile station to the closest point of $\Omega_{i}$ and returns from the farthest point, repeating this pattern for the subsequent reference sets. The extended generalized Fermat-Torricelli problem (5.3) can then be used to determine the optimal location of the mobile station.
Table 3 summarizes 24 cases with different parameter settings to evaluate the efficiency of the generalized SFT problem (extended variant) in accounting for wind effects. For simplicity, all cases are conducted on a small scale, with the slowest UAV assigned a nominal speed of 11(m/s); these values can be appropriately scaled to represent real-world scenarios. The results show that the generalized SFT problem reduces the transition time by aaproximately $66.8\%$. Furthermore, Cases 14 through 21 share identical settings except for the direction of the wind vector. Fig. 5(a) illustrates these results, confirming the ability of the generalized SFT formulation to address wind effects across varying directions.
7 Conclusion
After a disaster such as an earthquake, communication networks play a crucial role in supporting rescue operations. However, these networks are often partially or completely destroyed. In such scenarios, unmanned aerial vehicles (UAVs) can serve as valuable tools, offering large-scale WiFi coverage or Delay-Tolerant Networks (DTNs). To enable this, a fleet of UAVs can be deployed from a mobile station, such as a truck. Given the operational constraints of UAVs—such as battery limitation and the influence of external forces like wind—it becomes essential to determine the optimal location for the mobile station that minimizes both time and energy consumption. This challenge can be framed as a single facility location problem, which connects closely to two well-known problems in mathematical literature: the generalized Sylvester and Fermat–Torricelli problems.
The classical Sylvester and Fermat–Torricelli problems have long served as fundamental models in geometry and optimization. In this paper, we extended these problems into a more general framework that can include two types of gauge functions with several dynamic sets in $\mathbb{R}^{q}$, capturing the complexities of real-world facility location challenges. We introduced the generalized Sylvester–Fermat–Torricelli (SFT) problem, a novel framework for modeling single-facility location tasks in the presence of heterogeneous vehicle speeds, external forces like wind, and multiple distance norms (e.g., $\ell_{1}$ and $\ell_{2}$).
Figure 5: (a) Effect of wind direction on transition time at approximately similar wind speeds. The gray line shows transition time without considering wind, while the yellow line reflects results from the generalized SFT model accounting for wind. Results correspond to Cases 14–21 in Table 3. (b) Impact of wind vector norm (speed) on transition time. Results correspond to Cases 1–7 in Table 2
a) Refer to caption   b) Refer to caption
Several cases were studied involving multiple UAVs, and we demonstrated how the generalized SFT problem effectively addresses fleet heterogeneity and environmental conditions. Our results showed that this approach can reduce wasted operational time by up to 84%, highlighting its practical efficiency. Overall, our findings demonstrate the versatility of the generalized SFT problem in modeling practical optimization scenarios. Moreover, our proposed model can serve as a foundational framework for modern facility location problems—particularly in UAV applications—offering a basis for further development in future studies. Future research may explore further generalizations, including stochastic elements and multi-objective formulations, to broaden the scope and applicability of this foundational model. Also, other scenarios—such as multi-facility location problems and cases involving the $\ell_{1}$ norm—can be considered in future studies. In addition, integrating the generalized SFT problem with path planning algorithms can be investigated for last-mile delivery applications to minimize time, energy consumption, and operational costs.
Acknowledgments
This research is supported by the National Science Foundation via CMMI Grant 1944068.
8 Appendix: Required Lemmas, Propositons, and Tools
In this section, after some important definitions, we present the necessary propositions and lemmas that support the proof construction of the main propositions and the theorem discussed in the previous sections.
First, we define the affine hull of a set. Let $\Omega$ be a subset of $\mathbb{R}^{q}$, its affine hull is:
$\textnormal{aff}\ \Omega=\{\sum_{i=1}^{n}\lambda_{i}\omega_{i}\ |\ \sum_{i=1}^{n}\lambda_{i}=1,\lambda_{i}\in\mathbb{R},\ \omega_{i}\in\Omega,\ n\in\mathbb{N}\}.$ (8.1)
Next, we provide the definition of the relative interior of a set. Give $\Omega\subset\mathbb{R}^{q}$, the relative interior ($\textnormal{ri}\ \Omega$) of $\Omega$ is:
$\textnormal{ri }\Omega=\{x\in\Omega\ |\ \exists\ \epsilon>0\ \textnormal{such that}\ \mathbb{B}(x;\epsilon)\cap\textnormal{aff }\Omega\subset\Omega\}.$ (8.2)
Last, let $F\subset\mathbb{R}^{q}$ be a compact convex set with $0\in\textnormal{int }F$. Then, for a point $x\in\mathbb{R}^{q}$ and $r\geq 0$, the generalized ball centered at $x$ with radius $r$ is defined as follows
$B_{F}(x;r):=x+rF.$ (8.3)
8.1 Some properties of the Set-based Minkowski Gauge function
At the beginning of this section, we provide some properties for the generalized projection, defined in (3.5). The following proposition shows that $\Pi_{F}(x;\Omega)$ is nonempty.
Proposition 8.1.
Let $F\subset\mathbb{R}^{q}$ be a nonempty compact convex set with $0\in\textnormal{int }F$, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty closed convex set. Then, we have $\Pi_{F}(x;\Omega)\neq\emptyset$ for $x\in\mathbb{R}^{q}$.
Proof.
Consider the sublevel set
$\mathcal{L}_{\lambda}(\rho_{F}):=\left\{\omega\in\Omega\;\middle|\;\rho_{F}(x-\omega)\leq\lambda\right\},\quad\lambda\in\mathbb{R}.$
By Proposition 3.3, the function $\rho_{F}$ is continuous, and therefore, considering the closeness of $\Omega$, it is straightforward to show that $\mathcal{L}_{\lambda}(\rho_{F})$ is a closed set.
Furthermore, since $\rho_{F}(x-\omega)\to\infty$ as $\omega\to\infty$, we conclude that $\mathcal{L}_{\lambda}(\rho_{F})$ is also bounded; hence, $\mathcal{L}_{\lambda}(\rho_{F})$ is compact in $\mathbb{R}^{q}$. Because $\rho_{F}$ is continuous and $\mathcal{L}_{\lambda}(\rho_{F})\subset\Omega$ is compact, the Weierstrass Theorem guarantees that the infimum of $\rho_{F}(x-\omega)$ over $\mathcal{L}_{\lambda}(\rho_{F})$ is attained (which is equal to the infimum of $\rho_{F}(x-\omega)$ over $\Omega$). Hence, the solution set
$\Pi_{F}(x;\Omega):=\operatorname*{arg\,min}_{\omega\in\Omega}\rho_{F}(x-\omega)$
is nonempty.
The following proposition shows that $\Pi_{F}(x;\Omega)$ is a singleton when $F$ is strictly convex.
Proposition 8.2.
Let $F$ be a nonempty compact and strictly convex set with $0\in\textnormal{int}(F)$, and let $\Omega$ be a nonempty closed convex subset of $\mathbb{R}^{q}$, the generalized projection set $\Pi_{F}(x;\Omega)$ is a singleton.
Proof.
Based on proposition 8.1, we have $\Pi_{F}(x;\Omega)\neq\emptyset$. By contradiction suppose that $u,v\in\Pi_{F}(x;\Omega)$, then $\rho_{F}(x-u)=\rho_{F}(x-v)=r$, and therefore, $\frac{x-u}{r},\frac{x-v}{r}\in\textnormal{bd}F$. The function $\rho_{F}$ is subadditive and positively homogeneous (Theorem 6.14 in [22] provides its proof), so we have:
$\rho_{F}(x-\frac{u+v}{2})=\rho_{F}(\frac{1}{2}(x-u)+\frac{1}{2}(x-v))\leq\frac{1}{2}\rho_{F}(x-u)+\frac{1}{2}\rho_{F}(x-v)=r.$
Therefore, $\frac{u+v}{2}\in\textnormal{bd}F$ which is contradiction with strictly convex of $F$.
The following proposition shows that, under certain assumptions, the point belonging to the generalized projection lies on the boundary of the reference set.
Proposition 8.3.
Let F be a nonempty compact and convex subset of $\mathbb{R}^{q}$ such that $0\in int(F)$, and let $\Omega$ be a nonempty closed convex set in $\mathbb{R}^{q}$ with $\bar{x}\not\in\Omega$. If $\bar{\omega}\in\Pi_{F}(\bar{x};\Omega)$, then we get $\bar{\omega}\in\textnormal{bd }\Omega$.
Proof.
By contradiction, suppose $\bar{\omega}\in\textnormal{int}(\Omega)$, then, there exists $\epsilon>0$ such that $\mathbb{B}(\bar{\omega};\epsilon)\subset\Omega$. suppose $y=\bar{\omega}+\epsilon\frac{\bar{x}-\bar{\omega}}{\|\bar{x}-\bar{\omega}\|}$; therefore, we have:
$\rho_{F}(\bar{x}-y)=\rho_{F}(\bar{x}-\bar{\omega}-\epsilon\frac{\bar{x}-\bar{\omega}}{\|\bar{x}-\bar{\omega}\|})=\rho_{F}((\bar{x}-\bar{\omega})(1-\frac{\epsilon}{\|\bar{x}-\bar{\omega}\|}))=(1-\frac{\epsilon}{\|\bar{x}-\bar{\omega}\|})\rho_{F}(\bar{x}-\bar{\omega}).$
Since $(1-\frac{\epsilon}{\|\bar{x}-\bar{\omega}\|})\rho_{F}(\bar{x}-\bar{\omega})<\rho_{F}(\bar{x}-\bar{\omega})$, this is contradictin with the definition of generalized projection. Therefore, $\bar{\omega}$ should be a boundary point. ∎
The following proposition examines the conditions under which the subadditivity inequality of $\rho_{F}$ becomes an equality.
Proposition 8.4.
Let $F$ be a compact, strictly convex subset of $\mathbb{R}^{q}$ with $0\in\textnormal{int}F$. Then, the following equality holds
$\rho_{F}(x+y)=\rho_{F}(x)+\rho_{F}(y)\ \ \textnormal{with}\ \ x,y\neq 0$ (8.4)
if and only if $x=\lambda y$ for some $\lambda>0$.
Proof.
The proof can be found in Proposition 8.13 of [22].
The following lemma is a useful tool.
Lemma 8.5.
Let F be a nonempty closed convex set with $0\in\textnormal{int}F$, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty set. Then, we have $\Omega+rF=\Omega_{r}$ for $r>0$.
Proof.
For right inclusion, $x\in\Omega+rF$, so, $\rho_{F}^{\Omega}(x)\leq r$ and $x\in\Omega_{r}$. For the left inclusion, pick $x\in\Omega_{r}$ and we have $\rho_{F}^{\Omega}(x)\leq r$. Suppose $t\leq r$ and $\rho_{F}^{\Omega}(x)=t$, so $x\in\Omega+tF$. Since $F$ is convex and has 0 as the interior point, we have:
$\frac{t}{r}F\subset F\Rightarrow\Omega+tF\subset\Omega+rF.$
Therefore, $x\in\Omega+rF$.
The two following lemmas provide useful properties of the $\rho_{F}^{\Omega}$ function.
Lemma 8.6.
Let F be a nonempty closed convex subset in $\mathbb{R}^{q}$, let $\Omega$ be a nonempty set in $\mathbb{R}^{q}$, and let $r>0$. Suppose $x\not\in\Omega_{r}$ and $\rho_{F}^{\Omega}(x)<\infty$. Then we have
$\rho_{F}^{\Omega}(x)=\rho_{F}^{\Omega_{r}}(x)+r.$
Proof.
Given $\epsilon>0$, and let $\rho_{F}^{\Omega_{r}}(x)\leq t<\rho_{F}^{\Omega_{r}}(x)+\epsilon$ for some $\epsilon>0$. Then there exist $u\in\Omega_{r}$ and $f\in F$ such that $u+tf=x$. Since $\rho_{F}^{\Omega}(u)\leq r$, there exist $\alpha>0$, and $\rho_{F}^{\Omega}(u)\leq t^{\prime}<\rho_{F}^{\Omega}(u)+\alpha$ and $\omega\in\Omega$, and $f^{\prime}\in F$ such that $\omega+t^{\prime}f^{\prime}=u$. Substituting into the expression for $x$, we obtain:
$x=\omega+t^{\prime}f^{\prime}+tf.$
By the convexity of $F$, we have $t^{\prime}f^{\prime}+tf\in t^{\prime}F+tF=(t^{\prime}+t)F$, so
$x\in\omega+(t^{\prime}+t)F.$
This implies $\rho_{F}^{\Omega}(x)\leq t^{\prime}+t<\rho_{F}^{\Omega}(u)+\rho_{F}^{\Omega_{r}}(x)+\alpha+\epsilon$. Taking the limit as $\epsilon\to 0$ and $\alpha\to 0$, we conclude:
$\rho_{F}^{\Omega}(x)\leq\rho_{F}^{\Omega}(u)+\rho_{F}^{\Omega_{r}}(x)\leq\rho_{F}^{\Omega_{r}}(x)+r.$
For inverse inequality, suppose $\rho_{F}^{\Omega}(x)=t$, so $x\in\Omega+tF$. Since $x\notin\Omega_{r}$, we must have $r<t$. Then, convexity of $F$ gives us tF=(r+tr)F=rF+(tr)FtF=(r+t-r)F=rF+(t-r)F, and we have:
$x\in\Omega+rF+(t-r)F.$
Using Lemma 8.5 and the convexity of $F$, it follows that:
$x\in\Omega_{r}+(t-r)F,$
which implies $\rho_{F}^{\Omega_{r}}(x)\leq t-r=\rho_{F}^{\Omega}(x)-r$. Therefore,
$\rho_{F}^{\Omega_{r}}(x)+r\leq\rho_{F}^{\Omega}(x).$
Lemma 8.7.
Let F be a nonempty closed convex set in $\mathbb{R}^{q}$, let $\Omega$ be a nonempty set in $\mathbb{R}^{q}$, and let $t\geq 0$ and $f\in F$ and $x\in\textnormal{dom}\ \rho_{F}^{\Omega}$. Then,
$\rho_{F}^{\Omega}(x+tf)\leq\rho_{F}^{\Omega}(x)+t.$
Proof.
Given $\epsilon>0$, choose $h\geq 0$ such that $\rho_{F}^{\Omega}(x)\leq h<\rho_{F}^{\Omega}(x)+\epsilon;$ then, we have $x\in\Omega+hF$.
Since $tf\in tF$, it follows that
$x+tf\in\Omega+hF+tF=\Omega+(h+t)F,$
where the equality holds due to the convexity of $F$.
Hence, $\rho_{F}^{\Omega}(x+tf)\leq h+t<\rho_{F}^{\Omega}(x)+\epsilon+t.$ Taking the limit as $\epsilon\to 0$, we obtain
$\rho_{F}^{\Omega}(x+tf)\leq\rho_{F}^{\Omega}(x)+t.$
The following proposition states an important property of the set-based Minkowski gauge function.
Proposition 8.8.
Suppose $F\subset\mathbb{R}^{q}$ be a nonempty compact convex set with $0\in\textnormal{int}(F)$, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty closed convex set. Then, if $\rho_{F}^{\Omega}(\bar{x})=t>0$, then $\bar{x}=\bar{\omega}+tf$ where $\bar{\omega}\in\textnormal{bd}\Omega$ and $f\in\textnormal{bd}F$.
Proof.
By the definition of the generalized projection (3.5), we know that
$\rho_{F}^{\Omega}(\bar{x})=\rho_{F}(\bar{x}-\bar{\omega})=t\quad\textnormal{for}\quad\bar{x}\notin\Omega,$
where $\bar{\omega}\in\Pi_{F}(\bar{x};\Omega)$. By Proposition 8.3, we conclude that $\bar{\omega}\in\textnormal{bd}(\Omega)$. Therefore, we can write $\bar{x}-\bar{\omega}\in tF$, and thus, $\bar{x}\in\bar{\omega}+tF$. Consequently, there exists $f\in F$ such that $\bar{x}=\bar{\omega}+tf$.
Now, suppose $f\in\textnormal{int}(F)$. Then, there exists $\epsilon>0$ such that $\mathbb{B}(f;\epsilon)\subset F$. Let
$v=\frac{\bar{x}-\bar{\omega}}{\|\bar{x}-\bar{\omega}\|}.$
Since $f+\epsilon v\in F$, it follows that $t(f+\epsilon v)\in tF$. Expanding this, we have
$tf+t\epsilon v\in tF\Rightarrow\bar{x}-\bar{\omega}+\frac{t\epsilon(\bar{x}-\bar{\omega})}{\|\bar{x}-\bar{\omega}\|}=(\bar{x}-\bar{\omega})\left(1+\frac{t\epsilon}{\|\bar{x}-\bar{\omega}\|}\right)\in tF.$
Thus, it follows that
$\bar{x}-\bar{\omega}\in\frac{t}{1+\frac{t\epsilon}{\|\bar{x}-\bar{\omega}\|}}F.$
Since $t>0$ and $\epsilon>0$, let
$t^{\prime}=\frac{t}{1+\frac{t\epsilon}{\|\bar{x}-\bar{\omega}\|}}<t.$
This leads to a contradiction, as we have found $t^{\prime}$ such that $\bar{x}-\bar{\omega}\in t^{\prime}F$ and $t^{\prime}<\rho_{F}^{\Omega}(\bar{x})$. Therefore, $f\in\textnormal{bd}(F)$.
In the following proposition, we provide a relationship to calculate the minimal time to reach from point $x\in\mathbb{R}^{q}$ to the closest point in a set $\Omega\subseteq\mathbb{R}^{q}$.
Proposition 8.9.
Let $F\subseteq\mathbb{R}^{q}$ be a nonempty closed convex set with $0\in\textnormal{int}F$, and let $\Omega$ be a nonempty set. Then we have the following equality:
$\inf_{\omega\in\Omega}\rho_{F}^{\{\bar{x}\}}(\omega)=\rho_{-F}^{\Omega}(\bar{x})$
Proof.
For $\epsilon>0$, choose $t>0$, $\rho_{-F}^{\Omega}(\bar{x})\leq t<\rho_{-F}^{\Omega}(\bar{x})+\epsilon$. We can find $\omega\in\Omega$ and $f\in F$ such that $\bar{x}=\omega+t(-f)$. Therefore, $\omega=\bar{x}+tf$, and we have:
$\inf_{\omega\in\Omega}\rho_{F}^{\{\bar{x}\}}(\omega)\leq\rho_{F}^{\{\bar{x}\}}(\omega)\leq t<\rho_{-F}^{\Omega}(\bar{x})+\epsilon.$
Hence, it is obtained that $\inf_{\omega\in\Omega}\rho_{F}^{\{\bar{x}\}}(\omega)\leq\rho_{-F}^{\Omega}(\bar{x})$, as $\epsilon\rightarrow 0$.
For the inverse inequality, take $\alpha>0$, and choose $\inf_{\omega\in\Omega}\rho_{F}^{\{\bar{x}\}}(\omega)\leq d<\inf_{\omega\in\Omega}\rho_{F}^{\{\bar{x}\}}(\omega)+\alpha$. Then, there exists $\omega\in\Omega$ and $f\in F$ such that $\omega=\bar{x}+df$. Therefore, $\bar{x}=\omega+d(-f)$, and we have:
$\rho_{-F}^{\Omega}(\bar{x})\leq d<\inf_{\omega\in\Omega}\rho_{F}^{\{\bar{x}\}}(\omega)+\alpha.$
As $\alpha\rightarrow 0$, we conclude that $\rho_{-F}^{\Omega}(\bar{x})\leq\inf_{\omega\in\Omega}\rho_{F}^{\{\bar{x}\}}(\omega)$.
In the following proposition, we have an equality to calculate the minimal time to move from a point $\bar{x}\in\mathbb{R}^{q}$ to the farthest point of a set $\Omega\subseteq\mathbb{R}^{q}$.
Proposition 8.10.
Let $F\subseteq\mathbb{R}^{q}$ be a nonempty closed convex set with $0\in\textnormal{int}F$, and let $\Omega\subseteq\mathbb{R}^{q}$ be a nonempty bounded set. Then we have the following equality:
$\sup_{\omega\in\Omega}\rho_{F}^{\{\bar{x}\}}(\omega)=\bar{\rho}_{-F}^{\Omega}(\bar{x})$
Proof.
For $\epsilon>0$, choose $t\geq 0$ such that
$\bar{\rho}_{-F}^{\Omega}(\bar{x})\leq t<\bar{\rho}_{-F}^{\Omega}(\bar{x})+\epsilon.$
Then $\bar{x}-\Omega\subseteq tF$, or equivalently, for all $\omega\in\Omega$, we have $\bar{x}-\omega\in t(-F)$, which is the same as $\omega-\bar{x}\in tF$. This implies that
$\rho_{F}(\omega-\bar{x})\leq t\quad\text{for all }\omega\in\Omega.$
Using Proposition 3.1, we know that $\rho_{F}^{\{\bar{x}\}}(\omega)=\rho_{F}(\omega-\bar{x})$. Therefore,
$\sup_{\omega\in\Omega}\rho_{F}(\omega-\bar{x})=\sup_{\omega\in\Omega}\rho_{F}^{\{\bar{x}\}}(\omega)\leq t<\bar{\rho}_{-F}^{\Omega}(\bar{x})+\epsilon.$
Letting $\epsilon\to 0$, we conclude that
$\sup_{\omega\in\Omega}\rho_{F}^{\{\bar{x}\}}(\omega)\leq\bar{\rho}_{-F}^{\Omega}(\bar{x}).$
For the reverse inequality, let $d\geq 0$ such that
$\sup_{\omega\in\Omega}\rho_{F}(\omega-\bar{x})=\sup_{\omega\in\Omega}\rho_{F}^{\{\bar{x}\}}(\omega)=d.$
Then, for all $\omega\in\Omega$, we have $\rho_{F}(\omega-\bar{x})\leq d$, which implies $\omega-\bar{x}\in dF$. Hence, $\Omega-\bar{x}\subseteq dF$, or equivalently, $\bar{x}-\Omega\subseteq d(-F)$. Therefore,
$\bar{\rho}_{-F}^{\Omega}(\bar{x})\leq d.$
In the following proposition, we provide another formula for the subdifferential of the set-based Minkowski gauge function.
Proposition 8.11.
Let $F\subset\mathbb{R}^{q}$ be a nonempty compact convex set with $0\in\textnormal{int}F$, and let $\Omega$ be a nonempty closed convex set. For $\bar{x}\notin\Omega$ and $\bar{\omega}\in\Pi_{F}(\bar{x};\Omega)$, take $\rho_{F}^{\Omega}(\bar{x})=r$; then we have:
$\partial\rho_{F}^{\Omega}(\bar{x})=\partial\rho_{F}(\bar{x}-\bar{\omega})\cap N(\bar{x};\Omega_{r}),$ (8.5)
and
$\partial\rho_{F}^{\Omega}(\bar{x})=\partial\rho_{F}(\bar{x}-\bar{\omega})\cap N(\bar{\omega};\Omega).$ (8.6)
Proof.
Proposition 8.1 ensures that $\Pi_{F}(\bar{x};\Omega)$ is nonempty. Also, Proposition 3.6 gives $\partial\rho_{F}^{\Omega}(\bar{x})\subseteq\partial\rho_{F}(\bar{x}-\bar{\omega})$, and Proposition 3.5 gives $\partial\rho_{F}^{\Omega}(\bar{x})\subseteq N(\bar{x};\Omega_{r})$. Hence,
$\partial\rho_{F}^{\Omega}(\bar{x})\subseteq\partial\rho_{F}(\bar{x}-\bar{\omega})\cap N(\bar{x};\Omega_{r}).$
For the reverse inclusion $\supseteq$, it suffices to show that $\partial\rho_{F}(\bar{x}-\bar{\omega})\subseteq\Upsilon$, where $\Upsilon$ is defined in Proposition 3.5. Note that $\rho_{F}^{\{0\}}(x)=\rho_{F}(x)$, which implies $\partial\rho_{F}^{\{0\}}(x)=\partial\rho_{F}(x)$ for all $x\in\textnormal{dom }\rho_{F}$. In particular, at the point $\bar{x}-\bar{\omega}$ where $\rho_{F}(\bar{x}-\bar{\omega})=\rho_{F}^{\{0\}}(\bar{x}-\bar{\omega})=r$, let $v\in\partial\rho_{F}(\bar{x}-\bar{\omega})$. Then $v\in\partial\rho_{F}^{\{0\}}(\bar{x}-\bar{\omega})$, and by Proposition 3.5, we have
$v\in N(\bar{x}-\bar{\omega};\{0\}+rF)\cap\Upsilon,$
where $\{0\}+rF$ denotes the $r$-enlargement of the set $\{0\}$, as described in Lemma 8.5. Hence, $v\in\Upsilon$ and therefore, $\partial\rho_{F}(\bar{x}-\bar{\omega})\cap N(\bar{x};\Omega_{r})\subset N(\bar{x};\Omega_{r})\cap\Upsilon$. Using Proposition 3.5, $\partial\rho_{F}^{\Omega}(\bar{x})=N(\bar{x};\Omega_{r})\cap\Upsilon$, so $\partial\rho_{F}(\bar{x}-\bar{\omega})\cap N(\bar{x};\Omega_{r})\subseteq\partial\rho_{F}^{\Omega}(\bar{x})$. Therefore, we have:
$\partial\rho_{F}^{\Omega}(\bar{x})=\partial\rho_{F}(\bar{x}-\bar{\omega})\cap N(\bar{x};\Omega_{r})$
For equation (8.6), we have to show that $\partial\rho_{F}^{\Omega}(\bar{x})\subseteq N(\bar{\omega};\Omega)$. Take $v\in\partial\rho_{F}^{\Omega}(\bar{x})$ and fix $\bar{\omega}\in\Pi_{F}(\bar{x};\Omega)$, so that $\bar{x}-\bar{\omega}\in rF$. Then find $\bar{f}\in F$ such that $\bar{x}-\bar{\omega}=r\bar{f}$. For any $y\in G(\bar{x})$, where
$G(\bar{x}):=\left\{\omega+(\bar{x}-\bar{\omega})\,\middle|\,\omega\in\Omega,\,\bar{\omega}\in\Pi_{F}(\bar{x};\Omega)\right\}.$
Then, we have $y=\omega+r\bar{f}$ for some $\omega\in\Omega$. Now, we can write the subgradient inequality for $y$, noting that $\rho_{F}^{\Omega}(y)\leq r$. Therefore:
$\langle v,y-\bar{x}\rangle=\langle v,\omega-\bar{\omega}\rangle\leq\rho_{F}^{\Omega}(y)-\rho_{F}^{\Omega}(\bar{x})\leq 0.$
Thus, $v\in N(\bar{\omega};\Omega)$, and also, $\partial\rho_{F}^{\Omega}(\bar{x})\subseteq N(\bar{\omega};\Omega)$. This gives us $\partial\rho_{F}^{\Omega}(\bar{x})\subseteq\partial\rho_{F}(\bar{x}-\bar{\omega})\cap N(\bar{\omega};\Omega)$ where $\partial\rho_{F}^{\Omega}(\bar{x})\subseteq\partial\rho_{F}(\bar{x}-\bar{\omega})$ is obtained by Proposition 3.6.
For the inverse inclusion ($\supseteq$), we need to show that
$\partial\rho_{F}(\bar{x}-\bar{\omega})\cap N(\bar{\omega};\Omega)\subseteq\partial\rho_{F}(\bar{x}-\bar{\omega})\cap N(\bar{x};\Omega_{r})=\partial\rho_{F}^{\Omega}(\bar{x}).$
It suffices to show that for each $v\in\partial\rho_{F}(\bar{x}-\bar{\omega})\cap N(\bar{\omega};\Omega)$, we have $v\in N(\bar{x};\Omega_{r})$. Pick any $x\in\Omega_{r}$, and we have $\Pi_{F}(x;\Omega)\neq\emptyset$ by Proposition 8.1. Then, find $f\in F$ and take $x=\omega+tf$ where $w\in\Pi_{F}(x;\Omega)$ and $t\leq r$.
Since $v\in\partial\rho_{F}(\bar{x}-\bar{\omega})=\partial\rho_{F}^{\{0\}}(\bar{x})$, we have $v\in\Upsilon$ and $\langle v,f\rangle\leq\sigma_{F}(v)\leq 1$. Then:
$\begin{split}\langle v,x-\bar{x}\rangle&=\langle v,\omega+tf-\bar{x}\rangle\\ &=t\langle v,f\rangle+\langle v,\omega-\bar{\omega}\rangle+\langle v,\bar{\omega}-\bar{x}\rangle\\ &\leq t+\langle v,\omega-\bar{\omega}\rangle+\langle v,\bar{\omega}-\bar{x}\rangle\\ &\leq r+\langle v,\omega-\bar{\omega}\rangle+\langle v,\bar{\omega}-\bar{x}\rangle.\end{split}$
Since $v\in N(\bar{\omega};\Omega)$, it follows that $\langle v,\omega-\bar{\omega}\rangle\leq 0$. Also, since $v\in\partial\rho_{F}(\bar{x}-\bar{\omega})$, we have:
$\langle v,\bar{\omega}-\bar{x}\rangle=\langle v,-(\bar{x}-\bar{\omega})\rangle\leq\rho_{F}(0)-\rho_{F}(\bar{x}-\bar{\omega})=-\rho_{F}^{\Omega}(\bar{x})=-r.$
Therefore, for all $x\in\Omega_{r}$, we get $\langle v,x-\bar{x}\rangle\leq 0$, which implies $v\in N(\bar{x};\Omega_{r})$. Finally, we conclude:
$\partial\rho_{F}(\bar{x}-\bar{\omega})\cap N(\bar{\omega};\Omega)\subseteq\partial\rho_{F}(\bar{x}-\bar{\omega})\cap N(\bar{x};\Omega_{r})=\partial\rho_{F}^{\Omega}(\bar{x}).$
The following lemma establishes a property of the function $\rho_{F}$ that is useful in the proof of Theorem 3.9.
Lemma 8.12.
Let $F\subset\mathbb{R}^{q}$ be a nonempty, closed, convex set such that $0\in\operatorname{int}F$. Then, for all $x\in\mathbb{R}^{q}$, we have
$\rho_{F}(x)=\rho_{-F}(-x),$
where $\rho_{F}$ denotes the Minkowski gauge associated with $F$.
Proof.
Let $t:=\rho_{F}(x)$. By the definition of the Minkowski gauge (2.5), we have $x\in tF$. This yields $x\in t(-F)$, which implies $\rho_{-F}(-x)\leq t=\rho_{F}(x)$.
Conversely, let $d:=\rho_{-F}(-x)$. Then $x\in d(-F)$, so $x\in dF$, which implies $\rho_{F}(x)\leq d=\rho_{-F}(-x)$.
Combining both inequalities gives the desired result:
$\rho_{F}(x)=\rho_{-F}(-x).$
Next, we examine the relationship between the set-based Minkowski gauge function $\rho_{F}^{\Omega}$ and the minimal time function $T^{F}_{\Omega}$, defined by:
$T^{F}_{\Omega}(x):=\inf\{t\geq 0\mid(x+tF)\cap\Omega\neq\emptyset\},\quad x\in\mathbb{R}^{q},$ (8.7)
where $F\subseteq\mathbb{R}^{q}$ is a nonempty closed convex set and $\Omega\subseteq\mathbb{R}^{q}$ is a nonempty set. The following lemma establishes this relationship.
Lemma 8.13.
Let $F\subseteq\mathbb{R}^{q}$ be a nonempty, closed and convex set with $0\in\textnormal{int}F$, and let $\Omega\subseteq\mathbb{R}^{q}$ be a nonempty set. Then, we have the equality $\rho_{-F}^{\Omega}(x)=T^{F}_{\Omega}(x)$ for $x\in\mathbb{R}^{q}$.
Proof.
Using Propositions 3.1 and 8.9, we have:
$\rho_{-F}^{\Omega}(x)=\inf_{\omega\in\Omega}\rho_{F}^{\{x\}}(\omega)=\inf_{\omega\in\Omega}\rho_{F}(\omega-x).$
Theorem 6.19 in [22] shows that $T_{\Omega}^{F}(x)=\inf_{\omega\in\Omega}\rho_{F}(\omega-x)$. Therefore, $\rho_{-F}^{\Omega}(x)=T^{F}_{\Omega}(x)$ is proved.
In the final part of this section, we examine the relationship between the MSMG function and the maximal time function $C_{\Omega}^{F}$, which has been studied in the context of the generalized Fermat–Torricelli and generalized Sylvester problems [29, 32], and is defined as follows:
$C_{\Omega}^{F}(x):=\inf\{\,t\geq 0\;|\;\Omega\subseteq x+tF\,\},\quad x\in\mathbb{R}^{q},$ (8.8)
where $F\subseteq\mathbb{R}^{q}$ denotes a nonempty closed bounded convex set and $\Omega\subseteq\mathbb{R}^{q}$ denotes a nonempty bounded set.
The lemma below formalizes the relationship between the MSMG function and the maximal time function.
Lemma 8.14.
Let $F\subseteq\mathbb{R}^{q}$ be a nonempty closed bounded convex set with $0\in\operatorname{int}F$, and let $\Omega\subseteq\mathbb{R}^{q}$ be a nonempty bounded set. Then it holds that $\bar{\rho}_{-F}^{\Omega}(x)=C_{\Omega}^{F}(x),\quad\forall\,x\in\mathbb{R}^{q}.$
Proof.
Using Proposition 3.8 we have $\bar{\rho}_{-F}^{\Omega}(x)=\sup_{\omega\in\Omega}\rho_{-F}(x-\omega)$. Now, we want to show that
$\sup_{\omega\in\Omega}\rho_{-F}(x-\omega)=\sup_{\omega\in\Omega}\rho_{F}(\omega-x).$
To do this, take $\epsilon>0$ and $t=\sup_{\omega\in\Omega}\rho_{-F}(x-\omega)-\epsilon$ such that $t>0$. There exists $\omega\in\Omega$ such that $x-\omega\notin t(-F)$ which gives $\omega-x\notin tF$. Therefore, $t<\rho_{F}(\omega-x)\leq\sup_{\omega\in\Omega}\rho_{F}(\omega-x)$, and we have:
$\sup_{\omega\in\Omega}\rho_{-F}(x-\omega)-\epsilon<\sup_{\omega\in\Omega}\rho_{F}(\omega-x)\Rightarrow\sup_{\omega\in\Omega}\rho_{-F}(x-\omega)\leq\sup_{\omega\in\Omega}\rho_{F}(\omega-x)\ \ \text{as}\ \ \epsilon\to 0.$
The reverse inequality $\sup_{\omega\in\Omega}\rho_{-F}(x-\omega)\geq\sup_{\omega\in\Omega}\rho_{F}(\omega-x)$ is obtained in the same manner, so we conclude that $\sup_{\omega\in\Omega}\rho_{-F}(x-\omega)=\sup_{\omega\in\Omega}\rho_{F}(\omega-x)$. By Proposition 1 in [23] we have:
$C_{\Omega}^{F}(x)=\sup_{\omega\in\Omega}\rho_{F}(\omega-x).$
Therefore, we obtain that
$\bar{\rho}_{-F}^{\Omega}(x)=\sup_{\omega\in\Omega}\rho_{-F}(x-\omega)=\sup_{\omega\in\Omega}\rho_{F}(\omega-x)=C_{\Omega}^{F}(x).$
8.2 Existence and uniqueness of the optimal solutions of the generalized Sylvester problem
In the first step, the sufficient conditions for the existence of the solutions of (2.9) are considered.
In this section, we focus on the generalized Sylvester problem.
Theorem 8.15.
In the setting of the generalized Sylvester problem (2.9), there exists an optimal solution if the set $\Omega_{0}\cap(\bigcap_{i=1}^{n}\Omega_{i}+\lambda F_{i})$ is bounded for all $\lambda\geq 0$. In particular, this condition is satisfied if at least one of the sets $\Omega_{i}$, for $i=0,\ldots,n$, is bounded.
Proof.
The sublevel set for $\lambda\geq 0$ is defined as
$\mathcal{L}_{\lambda}(Y)=\{x\in\Omega_{0}\mid Y(x)\leq\lambda\}.$
When $Y(x)\leq\lambda$, this implies that $\rho^{\Omega_{i}}_{F_{i}}(x)\leq\lambda$ for $i=1,\ldots,n$. For any closed convex set $\Omega_{i}$, we have $\rho^{\Omega_{i}}_{F_{i}}(x)\leq\lambda$ if and only if $x\in\Omega_{i}+\lambda F_{i}$. Thus, if $x\in\mathcal{L}_{\lambda}(Y)$, then $x\in\Omega_{0}$ and $x\in\Omega_{i}+\lambda F_{i}$ for all $i=1,\ldots,n$. This implies that
$x\in\Omega_{0}\cap\left(\bigcap_{i=1}^{n}\Omega_{i}+\lambda F_{i}\right),$
and consequently, we have
$\mathcal{L}_{\lambda}(Y)\subseteq\Omega_{0}\cap\left(\bigcap_{i=1}^{n}\Omega_{i}+\lambda F_{i}\right).$
The reverse inclusion $\supseteq$” is straightforward. Moreover, the boundedness of the set $\Omega_{0}\cap\left(\bigcap_{i=1}^{n}\Omega_{i}+\lambda F_{i}\right)$ implies the boundedness of $\mathcal{L}_{\lambda}(Y)$. The boundedness of $\mathcal{L}_{\lambda}(Y)$ for all $\lambda\geq 0$ guarantees the existence of a solution to the generalized Sylvester problem (2.9), as established by Theorem 7.9 in [22]. The last statement is obvious.
In the following proposition, we provide a property for the $\rho_{F}^{\Omega}$ function.
Proposition 8.16.
Let $F\subset\mathbb{R}^{q}$ be a nonempty compact strictly convex set with $0\in\textnormal{int}F$, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty, closed, and strictly convex set. Then, the convex function $\rho_{F}^{\Omega}$ is not constant on any line segment $a,b]\subset\mathbb{R}^{q}$ with $a\neq b$ and $a,b]\cap\Omega=\emptyset$.
Proof.
By contradiction, suppose $\rho_{F}^{\Omega}(a)=\rho_{F}^{\Omega}(b)=r$ and $u=\Pi_{F}(a;\Omega)$ and $v=\Pi_{F}(b;\Omega)$. If $u=v$, then $\rho_{F}^{\Omega}(a)=\rho_{F}(a-u)=\rho_{F}(b-u)=r$; hence, $\rho_{F}(\frac{a-u}{r})=\rho_{F}(\frac{b-u}{r})=1$. Therefore, $\frac{a-u}{r}$ and $\frac{b-u}{r}$ are in $\textnormal{bd }F$. By strictly convex of $F$, $\frac{1}{2}(\frac{a-u}{r})+\frac{1}{2}(\frac{b-u}{r})\in\textnormal{int}(F)$. Now, we have contrdiction by considering
$\rho_{F}^{\Omega}(\frac{a+b}{2})\leq\rho_{F}(\frac{a+b}{2}-u)<r.$
In the other case, when $u\neq v$, let $\lambda\in(0,1)$ such that
$r=\rho_{F}^{\Omega}(\lambda a+(1-\lambda)b)\leq\rho_{F}(\lambda a+(1-\lambda)b-(\lambda u+(1-\lambda)v))\leq\lambda\rho_{F}(a-u)+(1-\lambda)\rho_{F}(b-v)=r.$
Therefore, $\lambda u+(1-\lambda)v=\Pi_{F}(\lambda a+(1-\lambda)b;\Omega)$ which means, by considering Proposition 8.3, $\lambda u+(1-\lambda)v\in\textnormal{bd }\Omega$. This contradicts the strict convexity of $\Omega$.
In the second step, we consider the condition for the uniqueness of the optimal solution.
Theorem 8.17.
In the setting of the problem (2.9), suppose that the sets $F_{i}$ and $\Omega_{i}$ for $i=1,\ldots,n$ are strictly convex. Then the problem admits at most one optimal solution if and only if the intersection $\bigcap_{i=0}^{n}\Omega_{i}$ contains at most one point. If, in addition, at least one of the sets $\Omega_{i}$, for $i=0,1,2,\dots,n$, is bounded, the problem has a unique solution.
Proof.
Define $\mathcal{H}(x)=Y^{2}(x),\ x\in\mathbb{R}^{q}$ and consider the optimization problem
$\textnormal{minimize}\ \mathcal{H}(x)\ \ \textnormal{subject to}\ \ x\in\Omega_{0}$
We are going to show that if $\bigcap_{i=0}^{n}\Omega_{i}$ contains at most one element, then $\mathcal{H}$ is strictly convex on $\Omega_{0}$, and hence the problem (2.9) has at most one optimal solution. Suppose, $\mathcal{H}$ is not strictly convex on $\Omega_{0}$. By proposition 8.33 in [22], we find a line segment $a,b]\subset\Omega_{0}$ with $a\neq b$ and a real number $r\geq 0$ so that
$Y(x)=\max\{\rho^{\Omega_{i}}_{F_{i}}(x)\ |\ i=1,\ldots,n\}=r\ \textnormal{for all}\ x\in[a,b]$
It is clear that $r>0$, since otherwise $a,b]\subset\bigcap_{i=0}^{n}\Omega_{i}$, which contradict the assumption. Proposition 8.35 in [22] tells us that there exist a line segment $\alpha,\beta]\subset[a,b]$ with $\alpha\neq\beta$ and an index $i_{0}\in\{1,\ldots,n\}$ such that
$\rho^{\Omega_{i_{0}}}_{F_{i_{0}}}(x)=r\ \textnormal{for all}\ x\in[\alpha,\beta].$
which contradicts the Proposition 8.16. Conversely, let problem (2.9) has at most one solution. Suppose that the set $\bigcap_{i=0}^{n}\Omega_{i}$ contains more than one element. It is clear that $\bigcap_{i=0}^{n}\Omega_{i}$ is the solution set of (2.9) with the optimal value equal to zero. Thus this problem has more than one solution, a contradiction that completes the proof. The last statement can be derived using Theorem 8.15.
8.3 Existence and uniqueness of the optimal solutions of the generalized Fermat-Torricelli problem
Next, we consider the Fermat-Torricelli problem and provide the conditions under which it has a unique solution.
Proposition 8.18.
[…]
$T(x)=\sum_{i=1}^{m}\rho_{F_{i}}^{\Omega_{i}}(x).$
which is defined in (2.12) is strictly convex on the constraint set $\Omega_{0}$.
Proof.
It follows from Propositions 3.2 and 2.1 that $T(x)$ is a convex function. Suppose that $T(x)$ is not strictly convex. Then there exist $x,y\in\Omega_{0}$ with $x\neq y$, and $\lambda\in(0,1)$, such that
$T(\lambda x+(1-\lambda)y)=\lambda T(x)+(1-\lambda)T(y).$
Since each $\rho_{F_{i}}^{\Omega_{i}}$ is convex, it follows that
$\rho_{F_{i}}^{\Omega_{i}}(\lambda x+(1-\lambda)y)=\lambda\rho_{F_{i}}^{\Omega_{i}}(x)+(1-\lambda)\rho_{F_{i}}^{\Omega_{i}}(y)\quad\text{for all }\ i=1,2,\dots,n.$
By assumption, there exists an index $j$ such that $L(x,y)\cap\Omega_{j}=\emptyset$. Let $\omega\in\Pi_{F_{j}}(x;\Omega_{j})$ and $v\in\Pi_{F_{j}}(y;\Omega_{j})$. Then, we can write $\rho_{F_{j}}^{\Omega_{j}}(x)=\rho_{F_{j}}(x-\omega)$ and $\rho_{F_{j}}^{\Omega_{j}}(y)=\rho_{F_{j}}(y-v)$. Therefore, we obtain:
$\begin{split}\lambda\rho_{F_{j}}(x-\omega)+(1-\lambda)\rho_{F_{j}}(y-v)&=\lambda\rho_{F_{j}}^{\Omega_{j}}(x)+(1-\lambda)\rho_{F_{j}}^{\Omega_{j}}(y)\\ &=\rho_{F_{j}}^{\Omega_{j}}(\lambda x+(1-\lambda)y)\\ &\leq\rho_{F_{j}}(\lambda x+(1-\lambda)y-(\lambda\omega+(1-\lambda)v))\\ &\leq\lambda\rho_{F_{j}}(x-\omega)+(1-\lambda)\rho_{F_{j}}(y-v).\end{split}$
Therefore, $\lambda\omega+(1-\lambda)v\in\Pi_{F_{j}}(\lambda x+(1-\lambda)y;\Omega_{j})$. By Proposition 8.3, we know that $\omega\in\textnormal{bd}(\Omega_{j})$, $v\in\textnormal{bd}(\Omega_{j})$, and $\lambda\omega+(1-\lambda)v\in\textnormal{bd}(\Omega_{j})$. Since $\Omega_{j}$ is strictly convex, it follows that $\omega=v$. Furthermore, the function $\rho_{F}$ is positively homogeneous (as shown in Theorem 6.14 of [22]), so we can write:
$\rho_{F_{j}}(\lambda x+(1-\lambda)y-\omega)=\lambda\rho_{F_{j}}(x-\omega)+(1-\lambda)\rho_{F_{j}}(y-\omega)=\rho_{F_{j}}(\lambda(x-\omega))+\rho_{F_{j}}((1-\lambda)(y-\omega)).$
Since $x,y\notin\Omega_{j}$, it follows that $x-\omega,y-\omega\neq 0$. Also, by Proposition 8.4, we have $\alpha>0$ such that $\lambda(x-\omega)=\alpha(1-\lambda)(y-\omega)$, so
$x-\omega=\beta(y-\omega)\ \ \textnormal{for}\ \ \beta=\frac{\alpha(1-\lambda)}{\lambda}\neq 1$
[…]
The next theorem proves the uniqueness of the solutions for the generalized Fermat-Torricelli problem under certain assumptions.
Theorem 8.19.
In the setting of problem (2.11), the problem admits a unique optimal solution under the following assumptions:
  1. i)
    The sets $F_{i}$ and $\Omega_{i}$, $i=1,2,\dots,m$, are strictly convex.
  2. ii)
[…]
  3. iii)
    At least one of the sets $\Omega_{i}$ for $i=0,\dots,m$ is bounded.
Proof.
First, we prove the existence of the optimal solution. To do this, for $\alpha>0$, the sublevel set fot $T$ is defined as:
$\mathcal{L}_{\alpha}(T)=\{x\in\Omega_{0}\ |\ T(x)\leq\alpha\}.$
By defenition of the $\rho_{F}^{\Omega}$, we have:
$\mathcal{L}_{\alpha}\subseteq\{x\in\Omega_{0}\ |\ \rho_{F_{i}}^{\Omega_{i}}(x)\leq\alpha\ \ \textnormal{for all}\ \ i=1,2,\dots m\}\subseteq\Omega_{0}\cap(\bigcap_{i=1}^{m}(\Omega_{i}+\alpha F_{i}).$
Since one of the sets $\Omega_{i}$ is bounded, the sublevel set $\mathcal{L}_{\alpha}$ is also bounded. Therefore, the existence of a solution follows from Theorem 7.9 in [22].
The uniqueness of the optimal solution follows from the strict convexity of the function $T$, as established in Proposition 8.18.
8.4 Experiments information
In this section, we provide information about the reference sets and dynamic sets used in the experiments, as summarized in Table 4.
Table 4: Experimental Information and Parameters
No. Refrence sets Dynamic sets
Info 1 $\Omega_{0}=\mathbb{R}^{2}$,
$\Omega_{1}=\left\{(x,y)\in\mathbb{R}^{2}\mid\max\{|x-30|,|y-350|\}\leq 15\right\}$, $F_{1}=\left\{(f_{1},f_{2})\in\mathbb{R}^{2}\mid(f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 1\right\}$,
$\Omega_{2}=\left\{(x,y)\in\mathbb{R}^{2}\mid\max\{|x-210|,|y-10|\}\leq 15\right\}$, $F_{2}=\left\{(f_{1},f_{2})\in\mathbb{R}^{2}\mid(f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 1\right\}$,
$\Omega_{3}=\left\{(x,y)\in\mathbb{R}^{2}\mid\max\{|x-550|,|y-200|\}\leq 15\right\}$ $F_{3}=\left\{(f_{1},f_{2})\in\mathbb{R}^{2}\mid(f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 1\right\}$
Info 2 $\Omega_{0}=\mathbb{R}^{2}$,
$\Omega_{1}=\left\{(x,y)\in\mathbb{R}^{2}\mid\max\{|x-30|,|y-750|\}\leq 20\right\}$, $F_{1}=\left\{(f_{1},f_{2})\in\mathbb{R}^{2}\mid(f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 1\right\}$,
$\Omega_{2}=\left\{(x,y)\in\mathbb{R}^{2}\mid\max\{|x-550|,|y-10|\}\leq 10\right\}$, $F_{2}=\left\{(f_{1},f_{2})\in\mathbb{R}^{2}\mid(f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 1\right\}$,
$\Omega_{3}=\left\{(x,y)\in\mathbb{R}^{2}\mid\max\{|x-950|,|y-400|\}\leq 15\right\}$ $F_{3}=\left\{(f_{1},f_{2})\in\mathbb{R}^{2}\mid(f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 1\right\}$
Info 3 $\Omega_{0}=\mathbb{R}^{2}$,
$\Omega_{1}=\left\{(x,y)\in\mathbb{R}^{2}\mid\max\{|x-30|,|y-200|\}\leq 5\right\}$, $F_{1}=\left\{(f_{1},f_{2})\in\mathbb{R}^{2}\mid(f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 1\right\}$,
$\Omega_{2}=\left\{(x,y)\in\mathbb{R}^{2}\mid\max\{|x-150|,|y-10|\}\leq 5\right\}$, $F_{2}=\left\{(f_{1},f_{2})\in\mathbb{R}^{2}\mid(f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 1\right\}$,
$\Omega_{3}=\left\{(x,y)\in\mathbb{R}^{2}\mid\max\{|x-350|,|y-90|\}\leq 5\right\}$ $F_{3}=\left\{(f_{1},f_{2})\in\mathbb{R}^{2}\mid(f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 1\right\}$
Info 4 $\Omega_{0}=\mathbb{R}^{2}$,
$\Omega_{1}=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-30)^{2}}{10^{2}}+\frac{(y-350)^{2}}{10^{2}}\leq 1\},$ $F_{1}=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 2^{2}\},$
$\Omega_{2}=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-210)^{2}}{10^{2}}+\frac{(y-10)^{2}}{10^{2}}\leq 1\},$ $F_{2}=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 1^{2}\},$
$\Omega_{3}=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-550)^{2}}{10^{2}}+\frac{(y-200)^{2}}{10^{2}}\leq 1\}$ $F_{3}=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 3^{2}\}$
Info 5 $\Omega_{0}=\mathbb{R}^{2}$,
$\Omega_{1}=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-40)^{2}}{10^{2}}+\frac{(y-550)^{2}}{10^{2}}\leq 1\},$ $F_{1}=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 2^{2}\},$
$\Omega_{2}=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-410)^{2}}{10^{2}}+\frac{(y-20)^{2}}{10^{2}}\leq 1\},$ $F_{2}=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 1^{2}\},$
$\Omega_{3}=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-750)^{2}}{10^{2}}+\frac{(y-350)^{2}}{10^{2}}\leq 1\}$ $F_{3}=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 3^{2}\}$
Info 6 $\Omega_{0}=\mathbb{R}^{2}$,
$\Omega_{1}=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-40)^{2}}{10^{2}}+\frac{(y-550)^{2}}{10^{2}}\leq 1\},$ $F_{1}=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 2^{2}\},$
$\Omega_{2}=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-110)^{2}}{10^{2}}+\frac{(y-20)^{2}}{10^{2}}\leq 1\},$ $F_{2}=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 1^{2}\},$
$\Omega_{3}=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-650)^{2}}{10^{2}}+\frac{(y-150)^{2}}{10^{2}}\leq 1\},$ $F_{3}=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 3^{2}\},$
$\Omega_{4}=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-750)^{2}}{10^{2}}+\frac{(y-650)^{2}}{10^{2}}\leq 1\}$ $F_{4}=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 2^{2}\}$
Info 7 $\Omega_{0}=\mathbb{R}^{2}$,
$\Omega_{1}=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-10)^{2}}{10^{2}}+\frac{(y-2000)^{2}}{10^{2}}\leq 1\},$ $F_{1}=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 3^{2}\},$
$\Omega_{2}=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-3000)^{2}}{10^{2}}+\frac{(y-1700)^{2}}{10^{2}}\leq 1\},$ $F_{2}=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 1^{2}\},$
$\Omega_{3}=\{(x,y)\in\mathbb{R}^{2}\ |\ \frac{(x-1700)^{2}}{10^{2}}+\frac{(y-20)^{2}}{10^{2}}\leq 1\}$ $F_{3}=\{(f_{1},f_{2})\in\mathbb{R}^{2}\ |\ (f_{1}-s_{1})^{2}+(f_{2}-s_{2})^{2}\leq 2^{2}\}$
Note that the dynamic sets in this table have two parameters, $s_{1}$ and $s_{2}$, which are obtained from the vector $\mathcal{S}=(s_{1},s_{2})$ under the Wind Vector column in Tables 1, 2, and 3.